### v1

- DoFs
- degrees of freedom
- 1D
- one-dimensional
- 3D
- three-dimensional
- PCC
- piecewise constant curvature
- PD
- proportional-derivative
- FEM
- finite element method
- MOR
- model order reduction
- ROM
- reduced order model
- EoM
- equation of motion
- PDE
- partial differential equation
- ODE
- ordinary differential equation
- ODEs
- ordinary differential equations
- BVP
- boundary value problem
- CRT
- Cosserat rod theory
- KRT
- Kirchhoff rod theory
- EBRT
- Euler-Bernoulli rod theory
- PCS
- piecewise constant strain
- DER
- Discrete Elastic Rod
- CSM
- continuum soft manipulator
- PCM
- parallel continuum manipulator
- SMA
- shape memory alloy
- TCP
- twisted and coiled polymer
- CTR
- concentric tube robot
- ETR
- eccentric tube robot
- MB
- Model-Based
- MP
- Multi-Point
- TT
- trajectory tracking
- MPC
- Model Predictive Control
- SMC
- Sliding Mode Control
- ESC
- Energy-Shaping Control
- FL
- Feedback Linearization
- PH
- Pythagorean Hodograph
- ML
- Machine Learning
- RL
- Reinforcement Learning
- SL
- Supervised Learning
- ANN
- artificial neural network
- RNN
- recurrent neural network
- MLP
- multi layer perceptron
- DQL
- deep Q-learning
- DDPG
- deep deterministic policy gradient
- PPO
- proximal policy optimization
- DR
- domain randomization
- MIS
- minimally invasive surgery
Carlo Alessi & Egidio Falotico, Brain-Inspired Robotics Laboratory, The BioRobotics Institute, Scuola Superiore Sant’Anna, Pisa, Italy.
Rod models in continuum and soft robot control: a review
Abstract
Continuum and soft robots can positively impact diverse sectors, from biomedical applications to marine and space exploration, thanks to their potential to adaptively interact with unstructured environments. However, the complex mechanics exhibited by these robots pose diverse challenges in modeling and control. Reduced order continuum mechanical models based on rod theories have emerged as a promising framework, striking a balance between accurately capturing deformations of slender bodies and computational efficiency. This review paper explores rod-based models and control strategies for continuum and soft robots. In particular, it summarizes the mathematical background underlying the four main rod theories applied in soft robotics. Then, it categorizes the literature on rod models applied to continuum and soft robots based on deformation classes, actuation technology, or robot type. Finally, it reviews recent model-based and learning-based control strategies leveraging rod models. The comprehensive review includes a critical discussion of the trends, advantages, limits, and possible future developments of rod models. This paper could guide researchers intending to simulate and control new soft robots and provide feedback to the design and manufacturing community.
keywords:
Soft Robotics; Modeling, Control, and Learning for Soft Robots; Biologically-Inspired Robots1 Introduction
Continuum robots are defined as robots with distributed deformations throughout their structure, resulting in infinite degrees of freedom (DoFs). This characteristic gives rise to a hyper-redundant configuration space, allowing the robot tip to reach any point in the three-dimensional (3D) workspace with a virtually infinite number of configurations (Trivedi et al., 2008c). Soft robots, built with soft materials or deformable structures (Laschi et al., 2016), are continuum robots. Generally speaking, soft robots assume a variety of morphologies. In this paper, we refer to both classes with their partial overlap.
Continuum and soft robots deform elastically while safely interacting with unstructured environments thanks to the inherent compliance of soft materials (Laschi and Cianchetti, 2014; Laschi et al., 2016). Novel design and manufacturing technologies are accelerating the prototyping of highly dexterous robotic systems (Rus and Tolley, 2015; Wang et al., 2024). However, the modeling and control of continuum and soft robots are still challenging problems. Indeed, soft materials cause extreme hyper-redundancy and exhibit nonlinear properties such as hysteresis and stress softening. These challenges inspired the development of diverse modeling techniques (Armanini et al., 2023; Gilbert, 2021), each characterized by underlying assumptions, mathematical framework, and a trade-off between accuracy and computational cost. Soft robotics encompasses four model classes: data-driven, discrete, geometrical, and continuum mechanical (Fig. 1). Concurrently, innovative control strategies emerged to leverage these models (George Thuruthel et al., 2018; Della Santina et al., 2023).
Popular modeling methods are inherently data-driven (Laschi et al., 2023). Despite not requiring geometric and physical expertise, they rely on pseudo-random motion data to train artificial neural networks mapping actuation to task space (Chin et al., 2020). This robot-independent method effectively derived efficient forward models employed in learned policies for tracking (Centurelli et al., 2022; Piqué et al., 2022) and throwing (Bianchi et al., 2023). However, they are black-box, could suffer overfitting, and the burden of collecting interaction datasets limits their applicability.
Other approaches, like discrete material models, discretize continuous bodies a priori. For example, pseudo-rigid models represent them with a chain of rigid links connected by joints. While these provide satisfactory results for hyper-redundant arms (Venkiteswaran et al., 2019), they poorly approximate elastic structures. In addition, lumped-mass models employ an array of masses, springs, and dampers (Habibi et al., 2020). This modularity can model complex phenomena but could require intensive system identification.
Alternatively, geometrical models utilize geometrical assumptions on the deformed shape of the soft body. For instance, functional models describe the deformations using polynomials (Della Santina and Rus, 2019). In addition, piecewise constant curvature (PCC) models discretize continuous bodies in circular arcs with constant curvature (Xie et al., 2022). These models were effectively used within proportional-derivative (PD) task space controllers for various dynamic tasks (Fischer et al., 2022). Since most continuum robots have actuators parallel to the centerline producing constant-curvature deformations, PCC is viable for uniform and lightweight robots (Webster III and Jones, 2010). However, the suitability of these approximations could degrade when robots are subject to significant external forces or interactions with the environment.
Finally, continuum mechanical models characterize soft robots with continuous configuration spaces and define deformations in physical terms. They enable the simulation of physical interactions and the study of robot mechanics. The classical finite element method (FEM) accurately represents complex 3D geometries to the extent of high computational costs and involved mathematical formulations (Coevoet et al., 2017). Since such models would impede real-time control applications, their use is often limited to designing and simulating soft robotic components (Xavier et al., 2021) or as a benchmark. Nonetheless, recent model order reduction (MOR) techniques are making 3D mechanical models more affordable (Allard et al., 2007; Dubied et al., 2022). Since continuum and soft robots are slender, other methods employ reduced order models of elastic rods, one-dimensional (1D) objects with a dimension (the length) much larger than the others. Elastic rods effectively describe the deformations of slender bodies (i.e., stretching, shearing, bending, twisting), balancing computational efficiency and an accurate representation of complex mechanics. Therefore, rod theories are promising candidates for soft robotics modeling and control, which are the focus of this paper.
1.1 Article Contribution
In this study, we investigate the role of rod theories in the modeling, simulation, and control of continuum and soft robots. The main contributions are:
1.
review of the fundamentals of four major rod theories with a common formulation, which facilitates the comparison and highlights connections to other modeling strategies;
2.
review and analysis of rod models applied to continuum and soft robots, originally classified on deformation classes; and
3.
review and analysis of recent model-based and learning-based controllers leveraging rod models.
First, we summarize the mathematical fundamentals underlying the four main rod theories applied in soft robotics: Cosserat-Reissner, Kirchhoff-Love, Timoshenko, and Euler-Bernoulli (Sec. 2). The same notation facilitates finding similarities and differences between the rod theories and supports the analysis of deformation classes. Further, it highlights theoretical insights on spatial discretization (i.e., positions vs configurations), connections to other models (e.g., the polynomial curvature), and implications on the control problem. Also, we consider the general modeling of actuation-induced distributed loads with different actuator routings (Sec. 3). Then, we review and discuss rod models applied to continuum and soft robots classified by deformation, and eventually robot types or actuation systems (Sec. 4). The analyzed models could guide researchers intending to simulate new soft robots in selecting the most suited rod theory. In addition, it could provide feedback to the design and manufacturing community. Finally, we review and discuss recent model-based and learning-based controllers that leverage rod models (Sec. 5). Along the sections, we critically discuss trends, research gaps, advantages, and limitations of rod models. The paper concludes with an overview of the main remarks (Sec. 6).
1.2 Related Surveys
Soft robotics has been the subject of extensive surveys, reviews, and perspectives, which represent a wealth of knowledge and a growing interest in a thriving research field.
The seminal paper by Webster III and Jones (2010) reviews PCC kinematic models of continuum robots, emphasizing the robot-specific and robot-independent mappings and providing a perspective on further developments. Conversely, our work focuses on rod models, which generalize PCC approaches in case of significant external forces.
Recently, Armanini et al. (2023) presented a structured overview of all modeling approaches in soft robotics. They propose a classification in four model classes (continuum mechanics, geometrical, discrete material, and surrogate), providing the fundamental theoretical grounds and analyzing their applicability. However, the control problem is only touched marginally. The perspective by Mengaldo et al. (2022) provide a unified formulation for modeling the embodied intelligence of soft robots. It broadly overviews the theories explaining the practical significance of embodied intelligence. Conversely, this paper is a review/survey article specific to rod-based models and controllers. Another work reviews state-of-the-art methods for soft robot modeling and control, emphasizing the use of real-time FEM (Schegg and Duriez, 2022). Also, Zhang et al. (2022) broadly covers a long soft robotics pipeline: design, actuation methods, modeling, and control. It focuses on major technical concerns, providing a perspective on the development of continuum robots. Finally, Sadati et al. (2023) survey MOR techniques and ROM, clarifying their disentangled usage in soft robotics. Within this framework, they broadly overview modeling approaches, constitutive laws, and solution strategies. While they survey the theoretical aspects of modeling, we also provide insights into the experimental validation of rod models.
Regarding control, George Thuruthel et al. (2018) provide a broad overview of model-based and model-free control strategies for continuum soft manipulators, without concern for the modeling. Also, Wang and Chortos (2022) provide a broad review of actuation mechanisms and control strategies, including open-loop, closed-loop, and autonomous methods. They discuss emergent directions in the control-actuator interface, underactuation, and utilization of artificial intelligence. Della Santina et al. (2023) describe the control problem of CSM specifically for the model-based view. They present a unified formulation of soft robot dynamics independent of the modeling technique. Then, they discuss the shape control and tracking problems and explore open challenges (e.g., under-actuation, environment interactions, actuator dynamics, task space control) while surveying the literature. Conversely, we review and discuss rod-based controllers utilizing model-based strategies and deep Reinforcement Learning (RL).
In summary, this review on rod models complements related works with minor intersections.
2 Background on Rod Theories
An elastic rod is a quasi 1D body in which the length is much larger than the radius of its cross-section. Therefore, the only deformation type considered is along one axis, neglecting other deformations like cross-section deformation. In nature, there are many examples of slender elastic bodies (e.g., hairs, muscle fibers, DNA strands, flagella). In the artificial world, notable examples are continuum and soft robots. In most cases, characterizing the behavior of these robots as that of elastic rods provides a good approximation. The four main Rod Theories used in soft robotics are: Cosserat-Reissner, Kirchoff-Clebsch-Love, Euler-Bernoulli, Timoshenko-Ehrenfest. These models differ in the kinematic assumptions and strain modes considered. An elastic rod might exhibit (i) bending, (ii) torsion, (iii) stretching (i.e., elongation or compression), and (iv) shear. Fig. 2 shows the four rod theories and the modeled strain modes. In this work, we adopted frequently the Lie Algebra notation. The reader can find in Appendix B the definition of the employed operators.
2.1 Problem Statement
An Elastic Rod can be kinematically described through a time-variant 3D curve, usually called backbone. The length of this curve can be parameterized by a material curvilinear abscissa . Every point of the backbone can be written w.r.t. an inertial frame . Usually, in literature, the point of the backbone at is called base and at is called tip of the robot. Furthermore, every cross-section has associated geometrical and material quantities: the cross-section area , the second moment of area about the local axes ; the mass density of the rod , the Young Modulus of the material , and the shear Modulus , where is the Poisson coefficient.
2.2 Cosserat Rod Theory
The Cosserat rod theory (CRT) (Cosserat and Cosserat, 1896) describes all strain modes of the rod. This model is useful for describing CSM that interact with cluttered environments, in which shear and elongation deformations play a fundamental role. Each cross-section is associated with a reference system , whose axes are called directors. The relative roto-translation between and is expressed by the homogeneous matrix defined as
where is the rotation matrix that represents relative rotation. The main idea of the CRT is to allow every cross-section to freely rotate and translate relatively.
2.2.1 Kinematics
Let be the strain twist defined as
where are the angular strain modes (bending and torsion), are the linear strain modes (shear and elongation/compression), and is the spatial partial derivative.
Similarly, it is possible to define the velocity twist as
where are the angular and linear velocities of cross-section , and is the time partial derivative. The strain and velocity twists are both expressed in the local frame and represent the evolution of the rod over space and time. Thanks to the mixed partial derivatives equality, it is possible to derive a kinematic relation between the strain and velocity twist, such as
where denotes an integration variable instead of the material abscissa .
From a mathematical point of view, the main idea of the CRT can be well condensed in the Configuration Space, defined as
Definition (5) indicates the infinite DoFs of an elastic rod. The Configuration Space results in a functional space of curves in (Boyer et al., 2020). A Configuration Space with an infinite dimension is a unique feature of soft robots, making control very challenging.
2.2.2 Dynamics
To describe the equations of motion of an elastic rod, it is necessary to define three distributed wrenches along the length of the rod:
The Internal Forces Wrench expresses the internal load applied by the material, including elastic and damping effects. Assuming small strains, a linear viscoelastic constitutive model relates the strain field and the internal forces wrench .
(6) where is the stress-free strain twist, is the stiffness matrix and is the viscosity matrix.
The External Forces Wrench represents the distributed external load applied to the Rod. For instance, the effect of gravity can be computed as
(7) where is the gravity acceleration twist w.r.t. the inertial frame and is the cross-sectional inertia matrix.
The Actuation Forces Wrench expresses the internal active forces exerted by the actuators. Refer to Sec. 3 for more details.
After defining these quantities, it is possible to compute the Dynamics of an elastic rod using the Poincaré equations (Renda et al., 2018a)
The EoM (8) is a set of partial differential equations of a Cosserat Rod in the local frame. It is common also to find the same set of PDE in an explicit form and expressed in the inertial frame, such as
where denotes a variable expressed in the inertial frame, is the second moment of the area expressed in the inertial frame, and and are the contributions of the internal and actuation forces.
2.2.3 Strain Parameterization
The Configuration Space (5) and PDE (8) show the infinite DoFs of an elastic rod. From a control and numerical implementation perspective, it is convenient to find a method to discretize the continuous nature of the soft segment. Renda et al. (2020) and Boyer et al. (2020) proposed a Strain Parameterization, expressing the Kinematics and the Dynamics of the elastic rod in terms of strain twist . Let us again consider the definition of the strain twist (2), which is easily rewritten as . Assuming that the initial condition is known, the function can be univocally determined by the strain twist. The Configuration Space can be considered the Shape Space of the elastic rod, which is a functional space of the -parameterized curves in , such as
This functional space can be generated by an infinite-dimension basis matrix , such as
where and is a vector of generalized coordinates, with . The main idea of the discretization technique is to truncate the basis matrix to a finite number of columns, reducing the discretized Shape Space to . This approach allows the user to choose the degree of approximation or neglect certain strain modes. To solve the PDE, the proposed method uses the Magnus Expansion (Haier et al., 2006), resulting in a convenient Product of Exponentials, which is widely used in classical robotics (Murray et al., 1994)
where denotes the Magnus expansion of the strain twist. Thanks to that, it is possible to rewrite the Differential Kinematics with a well-known form, such as
where denotes the Soft Geometric Jacobian. Similarly, the Dynamics can be rewritten in a classical Lagrangian form
The EoM (14) can be solved using a standard time solver, such as Runge-Kutta or Explicit Euler, as implemented in the SoRoSim simulator (Mathew et al., 2022).
2.2.4 Discrete Elastic Rod
Another method to discretize the continuum nature of elastic rods was introduced in the pioneering work of Gazzola et al. (2018). First, the Authors derived the EoM (9), including an elongation/compression ratio defined as , where is the curvilinear abscissa in the rest configuration. This ratio is present because the length is parameterized using the curvilinear abscissa differently from (9). In the presence of axial stretching, the stretching ratio scales the geometrical quantities
where the bar sign indicates the geometric quantities in the rest configuration. The EoM can be finally written as
It is worth highlighting that (16) contains an additional contribution with respect to (8), which depends on the time derivative of stretching ratio . Furthermore, scaling the geometrical quantities partially relaxes the assumption of rigid cross-sections. To numerically resolve the EoM (16), they extended for a Cosserat rod the spatial discretization algorithm proposed in a previous study (Bergou et al., 2008), which discretizes the rod in a sequence of rigid segments connecting nodes. For each node, the EoM consider the interactions with the other nodes and external forces. In addition, it is possible to associate kinematic and dynamic quantities to each node and segment to solve (16). In the discrete domain, some quantities, such as curvature, must be expressed in an integrated form over the domain (Gazzola et al., 2018). This domain corresponds to the Voronoi region associated with the interior nodes ,
where is the length of the -th segment. Then, the discrete curvature and bending stiffness matrix can be written as
with . Finally, the discretized EoM of the rod can be rewritten in algorithmic form as follows:
where is the point-wise mass associated with the node, is a domain dilation factor, and are the discrete difference and the averaging operator defined in (Gazzola et al., 2018). In the first equation of (19), , and . In contrast, in the second equation, ranges , while still ranges from . This distinction arises from the definition of discrete curvature and bending stiffness within the interior nodes set.
For the time integration of the discretized EoM (19), the authors proposed the use of a Second-Order Position Verlet time integrator, which exhibits a good balance between numerical accuracy and computational cost (Gazzola et al., 2018). To ensure numerical stability, they proposed an empirical law to choose the integration frequency as . The discrete EoM (19) were implemented in the PyElastica simulator (Naughton et al., 2021).
2.3 Kirchhoff Rod Theory
The Kirchhoff rod theory (KRT) (Love, 1906) is a special case of CRT that considers an elastic rod unstretchable and unshearable. It is particularly suitable for CSM that bend around any axis and twist. Notably, it introduces the notion of a directed curve, assigning a specific reference system to each cross-section. The EoM of Kirchhoff rods can be computed by specializing the EoM of Cosserat, that is (8) or (9). In particular, the constraint of an unstretchable and unshearable rod can be written as
This constraint can be added in the EoM (16), resulting in
The linear internal force serves as a Lagrangian multiplier and it is a virtual internal force that constrains the rod from stretching or shearing. In the case of Strain Parameterization, it is sufficient to apply the constraint (20) in the strain twist
Therefore, the basis function matrix only generates the angular strain mode vector .
2.4 Euler-Bernoulli Rod Theory
The Euler-Bernoulli rod theory (EBRT) (Timoshenko, 1983) is one of the simplest rod theories in which the rod can only bend around one axis. It can be considered the 2D case of the KRT without the twist. Here, the assumption is that the slope angle of the rod is equal to the tangent angle of the backbone curve. Below, we report linear and nonlinear versions.
2.4.1 Linear Euler-Bernoulli
Let us consider the rod in the - plane. Unlike the previous theories, the backbone curve is described by the displacement from the -axis. Recalling the assumption of the EBRT, the slope angle of the beam can be written as
From the minimization of the strain energy, the EoM of the rod can be computed as
where is the distributed external moments around the -axis. The EoM (24) is linear, assuming that the constitutive law of the bending moment is
where is the -component of the internal moment . For control purposes, the linearity of (24) facilitates the use of efficient controllers from the classic control theory (Doroudchi et al., 2018).
2.4.2 Nonlinear Euler-Bernoulli
In the nonlinear EBRT (or Euler’s Elastica), the rod length is parameterized by the arclength . Recalling the Fundamental Theorem of the local theory of Curves in the 2D case, every regular curve can be determined by the curvature (do Carmo, 1976). From this concept, Della Santina and Rus (2019) reformulated Elastica in a classical robotic formulation. Let be the -component of the curvature twist . The Cartesian pose of every cross-section is uniquely determined by the curvature, i.e.
Furthermore, from (26), it is possible to derive the EoM following the Lagrangian approach.
Similarly to the Strain Parameterization, they propose to consider the curvature as an infinite sum of monomials in , i.e.
The geometrical meaning of the polynomial curvature is to constrain the shape of the backbone curve to be a Generalized Cornu Spiral, i.e., a curve with a polynomial curvature (Della Santina and Rus, 2019). It is worth highlighting that the (27) is equivalent to write (11) with a polynomial basis function, related only to the curvature (Caradonna et al., 2024). Defining as joint variables, it is possible to write the EoM as
where is transposed orientation jacobian for actuators and are the pure actuators’ torque.
2.5 Timoshenko Rod Theory
The Timoshenko-Ehrenfest rod theory (Timoshenko and Goodier, 1951) extends the EBRT, relaxing the unshearability constraint. Therefore, the equality between the tangent angle and the cross-sections’ angle is no longer valid (i.e., ). For this reason, the EoM of a Timoshenko rod considers the angle as an independent variable. Its dynamics can be written as
The standard formulation assumes a linear constitutive law with bending moment and shear force defined as
where is the Timoshenko shear coefficient, which depends on the cross-section geometry. We can see that the shear force contribution is proportional to the difference between the tangent angle and the angle of cross-section .
3 Modeling Actuation
The distinction between physical and robotic models lies in the inclusion of actuation. In soft robotics, actuation sources greatly differ from those used in rigid robots. In particular, the most used technologies are cables and fluidic chambers, which guarantee distributed active loads .
3.1 The Actuation Matrix
Consider a Cosserat rod with actuators. Let be the distance from the center of the cross-section and the -th actuator (Fig. 3). Each actuator applies an internal active wrench that depends on the actuator routing. We can express the internal active wrench applied by the actuator as
where is the tangent versor to the actuator path and is the magnitude of the i-th actuator. For example , can be
where is the cable tension, is the pneumatic pressure and is the internal cross-section area of the fluidic chamber. Moreover, the expression of the tangent versor is derived as
where is expressed in homogeneous coordinates and the operator extracts the first three rows of a homogeneous vector. The resultant of the contributions of actuators is
where in the internal active wrench and is the actuation matrix
The actuation matrix is crucial for designing control algorithms because it contains valuable information about the strain modes that can be excited by the actuators. From a control perspective, this information is related to the reachability of the system.
3.2 Actuators-Strain Mapping
From the definition of actuation matrix (35), it is possible to map the actuation magnitude vector to the distributed active load . Renda et al. (2020) propose a statics-based method to relate the actuators with excited strain modes. In particular, let be the statics of a Cosserat Rod as
assuming no external forces and the strain twist is discretized with the Strain Parameterization approach, i.e. (11). By invoking the D’Alembert Principle, it is possible to derive
Finally, substituting the constitutive relation (6) and the definition of actuation forces (34), we derive
From this equality, Renda et al. (2024) propose a trivial form of the static equation, choosing the implicit parametrization
Consequentially, the static equation degenerates in a trivial form, such as
The functional basis derived from the implicit parametrization (39) uses the information of actuation routing contained in , to provide the minimum set of functional basis. In particular, these functional bases describe which are the excited strain modes and which is the best function to describe them statically. Furthermore, in (35) considers also the geometrical and the material information of the robot, contained in . Finally, (39) can be a useful starting point to build the optimal functional basis matrix , avoiding redundant or inefficient shape functions, preserving accuracy and computational efficiency.
4 Rod Models for Continuum and Soft Robots
The rod models were classified based on combinations of the four deformation modes. Out of the fifteen possible combinations, only nine classes emerged. Fig. 4 shows the distribution of rod models over deformations classes. Within each deformation class, we divided the papers mainly by actuation, robot type, or topic. In each division, the articles follow a chronological or logical order. An advanced Scopus query aided the selection of the articles (see Appendix A). Fig. 5 shows how different robots might exhibit these deformations induced by actuators or the interaction with the environment. We survey rod models ordering the deformation classes from most to least popular (see Fig. 4).
We summarize the four major deformation classes (i.e., Bend, All Deformation, Bend & Twist, Bend & Stretch) respectively in Table 1-4 to facilitate the literature comprehension. Each table arbitrarily reports representative articles to distill the deformation class and highlight the versatility of rod models.
4.1 Bend
Bending is a prevalent deformation thanks to the slender morphology of continuum and soft robots (Fig. 5). Herein, we review the evolution of rod models undergoing bending through various mechanisms. Table 1 distills the deformation class reporting representative works.
4.1.1 Bend by Cables
The literature presents several solutions to predict the bending of soft robots. The CRT computed in real-time the kinematics of tendon-actuated robots subject to external loads , modeling the tendon actuation as a single point moment applied to the backbone where each tendon is attached, achieving position errors below on a Nitinol rod (Jones et al., 2009). Direct and inverse kinematic models were presented drawing inspiration from the structure of octopus arms (Giorelli et al., 2012). The direct kinematics exploited a Cosserat rod model. Conversely, the inverse kinematics computes the Jacobian matrix iteratively from the task space. The kinematic models overcome the slack problem as they work directly on the cable tensions instead of controlling the cable extension. Actuations validated the model on a prototype immersed in water. The challenge of predicting the bending angle of a planar manipulator driven by four wires and with triangular notches was addressed using a kinematic Timoshenko rod model (Wenlong et al., 2013). A two-node Timoshenko rod element modeled each triangular unit, accurately predicting the bending angles measured on the tensioned manipulator. Concerning planar grasping, it was proposed a dynamic Cosserat rod model for a tendon-driven continuum robotic finger (Dehghani and Moosavian, 2014b). The finger grasping is simulated as an external point force acting on the finger backbone , assuming that the objects do not slip. Also, a singularity-free version of the model equations is derived using the Taylor expansion. A preliminary experimental validation is conducted by attaching different weights on the tip of a superelastic NiTinol rod. Also, the EBRT was derived using the Lagrangian method to predict the shape of a multi-backbone continuum robot (He et al., 2013). The experimental validation obtained position errors lower than 4%L. Moreover, Gao et al. (2016) predicted manipulator deformations due to external forces, tendon forces, or friction. They provide a planar Cosserat rod model of dexterous continuum manipulators, achieving sub-millimeter tip error predictions in the experimental validation. A dynamic model of a cable-driven soft tentacle was developed to explore the relationship between the strain field and the rope length by formulating a rope-driven actuator model (Ma et al., 2021). A kinematic model based on CRT was presented. Then, a Newton-Euler inverse dynamics algorithm for an equivalent CSM computed the matrices of the Lagrangian inverse dynamics model. A computationally efficient analytical tip force estimator was proposed for tendon-driven catheters (Hooshiar et al., 2021). First, they used a Bezier-based shape interpolation assuming constant curvature. Then, they proposed an optimal problem to invert the system statics and compute the tip force . The experimental validation achieved a N mean absolute error at high frequency. Concerning the interaction with objects, a general CRT-based framework investigated the coupled dynamics of tendon-driven cooperative continuum robots interacting with flexible objects (Jalali and Janabi-Sharifi, 2021). The framework is validated experimentally using two cable-driven spring steel arms performing trajectories with a Nitinol object.
Workspace investigations are crucial for dexterous manipulation tasks. A study on estimating the workspace of CSM was conducted by developing a quasi-static Cosserat rod model and introducing a continuation algorithm for a tendon-driven CSM (Amehri et al., 2021). The algorithm maps the exterior and interior workspace boundaries, addressing challenges like actuator saturation, physical limitations, self-collision, or elastic instabilities. The study highlights that complex configurations of CSM can lead to isolated boundaries and voids. Validation through numerical simulations on planar and spatial configurations demonstrated the effectiveness of the proposal.
Other works focus on the stiffness analysis. A Cosserat rod model for a continuum manipulator with variable stiffness was used to compute the tip stiffness (Zhao et al., 2020a). Experimental results demonstrate the ability to adjust the tip stiffness in various directions with an increase of about ten times the minimum stiffness. Also, a novel cable-driven continuum robot with layer jamming sheaths, which control the stiffness in the transverse and axial directions, is designed, modeled, and validated (Fan et al., 2022). An Euler-Bernoulli rod and the Maxwell-Mohr method derived the effective robot stiffnesses, expressing the rod deflection as
where is the moment due to the unit load. Extensive static force-deflection tests are conducted with different layer jamming sheaths, vacuum pressures, and bending angles. A novel approach proposed additional cables to selectively stiffen planar cable-driven robots with limited coupling with the actuating cables (Molaei et al., 2022). This approach increases the control of shape and stiffness of continuum manipulators without significantly affecting the basic kinematics and addressing uncontrolled deformations like buckling. A planar Cosserat rod model based on (Rucker and Webster III, 2011) identified the design conditions to route the additional cables. The simulations show a contained coupling between the length changes of the actuating and stiffening cables. The proposal is validated experimentally on a prototype with two bending cables and two stiffening cables, demonstrating that the secondary tendons increase the stiffness by ten times. Stiffness is also investigated for a simple rod-shaped silicone soft robot modeled with a geometrically exact CRT (Grube and Seifried, 2022). The stiffness properties are investigated by applying horizontal pulling forces up to 3 N to the rod and robot free-ends. They analyzed experimentally a linear viscoelastic material model and the nonlinear Saint Venant-Kirchhoff material model. The analysis of the damping mechanism has shown that linear damping can be assumed for oscillations with small to medium amplitudes since nonlinear material laws improve the model accuracy only slightly. Therefore, nonlinear effects have to be considered for oscillations of higher amplitudes. Xiao et al. (2023) present an on-demand stiffness modulation approach that retains the intrinsic compliance of elastomeric bodies. They use a Nitinol tube as the backbone, which slides through a cable-driven soft body to achieve robot pose or stiffness modulation. They develop and validate a CRT model in different scenarios by varying and the contact force . Experimental results indicate average shape and tip errors below 1%L, while simulations demonstrate that backbone insertion enhances the workspace and reduces compliance by 57.7%. Their design was further validated by object manipulation and minimally invasive surgery (MIS) case studies.
A generalized kinetostatics CRT model with friction considered for multi-module and multi-segment continuum robots is proposed (Wang et al., 2023). The model is applied to a two-module rod-driven continuum robot with winding ropes to analyze its deformation and load characteristics. The experimental validation, with and without loads, achieves less than shape error.
4.1.2 Bend in Soft Pneumatic Actuators
Several works estimate the bending deflection in pneumatic soft robots. An Euler-Bernoulli rod described a bending pneumatic soft actuator (de Payrebrune and O’Reilly, 2016). Through experimental procedures with the robot clamped at one end, the curvature was expressed as a function of the applied pressure . The results show that the curvature field is a function of the applied pressure.
FEM simulations often provide a benchmark. Using FEM and experiments, a five-parameter constitutive relation for a rod-based pneumatic soft arm model was explored (De Payrebrune and O’Reilly, 2017). The constitutive law’s parameters are determined for pneumatic soft arms through FEM simulations under different conditions. Furthermore, various actuator sections are investigated (e.g., rectangular, semicircular). The parameterization is validated by comparing FEM with the rod model for various actuator sections, subject to tip loads . Another work introduced interoperable dynamic modeling for fluidic elastomeric soft fingers (Mbakop et al., 2021b). The model employs an Euler-Bernoulli framework with Pythagorean Hodograph (PH) curves reconstructing the kinematic shape from virtual control points. The shape of the soft finger is recovered as a function of the applied pressure and control points as
Here, is the height from the ventral part of the finger in the cross-section, and is the normalized curvilinear coordinate along the neutral axis of the soft finger. In this way, a relationship is established between control inputs and PH curve control points, enabling dynamic inversion. FEM simulations validated the PH-EBRT approach for single and two-segment soft fingers, while the experimental validation measures cartesian error and bending curvature on 3D-printed soft fingers of varying stiffness. Interestingly, EBRT successfully modeled a compliant robotic fish tail with fluidic actuators (Wu and Shi, 2022b). The model predicts actuator deflections and nonlinear strains. Nonlinear FEM validated the approach by comparing the results in terms of the first, second, and third obtained natural frequencies, showing good agreement. Finally, the bending of a beam-shaped soft actuator with an embedded fluidic network was analyzed by leveraging a dynamic Euler-Bernoulli rod with large deflections and nonlinear strains (Janizadeh Haji and Bamdad, 2023). Its accuracy is compared against nonlinear FEM and 4th Runge–Kutta. Examining multi-mode and nonlinear strain effects revealed instabilities in some frequencies for the nonlinear model. Geometric parameters influence actuator deformations, with increased channel distribution density enhancing deflection. Moreover, larger fluid channel diameter and actuator length require fewer inlet pressures for the same bending, while higher actuator width and height demand higher pressures. Conversely, Sadati et al. (2019) introduced a ROM and a discretized model based on EBRT with absolute states for the STIFF-FLOP. In addition, they improve the model in (Renda et al., 2018a) with an Euler-Bernoulli discretized model with relative states. An experimental comparison revealed that the ROM is more accurate and numerically robust, while absolute states provide higher computational efficiency. In contrast, relative states show higher sensitivity and computational burden. Finally, a lumped-mass model provided a benchmark with the highest computational cost among the considered models.
Bending structures also exhibit stiffness modulation and adaptation. For instance, CRT modeled a soft bending actuator with an added fabric layer, which changes the actuator stiffness, enhancing the bending and preventing longitudinal stretching (Nalkenani et al., 2021). The authors incorporate fabric effects as a boundary condition in the Cosserat model and validate experimentally across actuation frequencies and loading conditions, producing accurate behavior predictions. Also, it was proposed a nonlinear kinematic Cosserat rod model of a pneumatic surgical CSM with stiffness adaptation (Roshanfar et al., 2022). The model incorporates nonlinear material properties through the two-term Mooney-Rivlin hyperelastic constitutive law, whose strain energy density is
Here, and are the deviatoric strains invariants while and are material constants. The simulation and experimental results for tip displacement and stiffness show similar trends with a maximum error of 8.25%.
More recently, the work in (Namdar Ghalati et al., 2022) presented a method to obtain the static shape of a semicylindrical soft fiber-reinforced pneumatic bending actuator under external force constraints. The pressure-angle relationship in free motion is derived using the Neo-Hookean strain energy theory of hyper-elastic materials, i.e.,
Here, is the first invariant of the right Cauchy-Green deformation tensor and , , are the principal stretches (Odgen et al., 2013). They investigated the effects of external forces using the EBRT, which modeled the arm as a cantilever beam under external load.
4.1.3 Bending with Smart Actuators
Bending is achieved by several innovative actuation technologies. Inspired by the forward crawling motion of the Manduca sexta caterpillar, a dynamic Euler-Bernoulli rod model for a soft robot with shape memory alloy (SMA) actuators was introduced to analyze undulation mechanics (Daily-Diamond et al., 2017). The model employs a motion pattern based on the Witch of Agnesi function
Its shape is analogous to the probability density function of the Cauchy distribution. Lastly, experimental data on an SMA actuator’s flexural rigidity, curvature, and recorded caterpillar leg velocity enable studying ground contact forces and the energetics of successful forward locomotion. Differently, CRT solved direct and inverse kinematics of soft robots actuated by twisted and coiled polymers (Pawlowski and Zhao, 2018). Direct kinematics determines the manipulator shape based on TCP temperatures, and inverse kinematics uses an augmented Jacobian approach (Rucker and Webster, 2011). A subsequent work integrates a thermodynamic TCP model and a dynamic Cosserat rod model, with promising forward dynamics results (Pawlowski et al., 2018). Additionally, an Euler-Bernoulli rod model predicted the deflection of a micro soft robot made from a magnetic elastic composite material (Xiang et al., 2019). The 3D-printed robot, having a head and a tail for rolling and crawling, is controlled by two magnetic fields generated by a 3D Helmholtz coil. Experimental validation for rolling and crawling velocity under different input waveforms, amplitudes, and frequencies shows a linear relation of rolling speed to rotating magnetic field frequency. Moreover, crawling speed correlates positively with magnetic field amplitude and frequency, being faster with sinusoidal compared to triangular wave inputs. A recent work developed a dynamic model of a continuum robotic arm with a 3D-printed elastic backbone and orthogonal plates supporting distributed soft cubic actuators, composed of a novel temperature-responsive hydrogel, each with an embedded Joule heater (Doroudchi et al., 2020). Leveraging the CRT, they simplify the PDE into reduced ordinary differential equations by integrating semi-discretized distributed force and rotational inertia components. Assuming robot isotropy and homogeneity, they estimate the Young modulus and damping coefficients through vibrational tests. Experimental verification accurately predicts tip displacements with a maximum of error. Finally, a study proposes a model of CSM with magnetic actuation (Tariverdi et al., 2020). They use Lie group synchronous variational integrators to derive a geometrically exact model based on an extended dynamic Kirchhoff rod and Hamiltonian formulation. The model considers nonlinearities, magnetic field loads, gravity, and internal and external dissipation forces from friction. Experiments involve dynamic motions with metallic and polymeric soft rods. Polymer rods show higher errors due to fabrication uncertainties and nonlinear elastic properties. Increasing the number of nodes improves the system frequency accuracy.
4.1.4 Bending in Parallel Manipulators
A novel parallel continuum manipulator (PCM) constructed with soft panels and its shape modeling approach was presented (Li et al., 2018). They used the nonlinear EBRT to model the deformations of each soft panel. By solving a constrained optimization for the kinematics, they calculated the deformed shapes and obtained the relationship between the panels’ length and the pose of the moving platform. Experiments validated the model reaching good shape predictions.
Li et al. (2023) present a kinematic model of a planar PCM made of two independent flexible panels. The formulation unifies the geometric constraints, the load balance of the moving platform, and the panel’s statics. For each panel, the relation between internal moment and deflection is described by the EBRT via elliptic integrals, which account for inflection points with large deformations. Simulation results compared against FEM reveal that, under large deflections, the panels exhibit two inflection points close to both ends, which affect the distribution of the internal moment along the panel, (25). Stress checking confirms that the deformations are within the elastic region.
4.1.5 Bending in Tubular Robots
Bending is the prominent deformation of tubular robots aimed for MIS. Pattanshetti and Ryu (2019) employ the CRT to model a laser-patterned monolithic tube with varying structural properties along the length. The combination of CRT and a vision-based identification algorithm enables the characterization of the non-linearities and the bending stiffness along the rod with increasing strain. Another work adjusts the piecewise variable-strain (Renda et al., 2020; Boyer et al., 2020) to simulate multi-section concentric tube robots by including the tubes sliding motion (Renda et al., 2021). Rotation motions of the tubes are included as generalized coordinates instead of boundary kinematic conditions. This procedure led to a minimum set of closed-form algebraic equations solvable for both the shape variables, the actuation forces, and the torques. The resulting equations facilitate control, design optimization, and stability assessment. The work tackles one of the main limitations of the sliding structure of standard CTR, the spaghetti problem.
4.1.6 Bending in Simulated Locomotion
Bending is also exploited to investigate in simulation locomotion strategies (Fig. 5). Zhou et al. (2015b) describe a soft robotic limb attached to a mass with an Euler-Bernoulli rod to examine in quasi-static conditions the peristaltic locomotion controlling the intrinsic curvature on a rough horizontal surface. The paper also provides design recommendations for the optimal performance of soft robots with variable curvature and adhesion. Interestingly, they also studied grasping strategies using a gripper with two fingers attached to a palm modeled using two rods (Zhou et al., 2015a). Concerning snake-like locomotion, especially undulatory swimming, is examined by studying a planar inextensible elastic Cosserat rod with adjustable spontaneous curvature, providing an internal actuation mechanism resembling muscular action in a snake (Cicconofri and DeSimone, 2015). Equations of motion are derived for (a) the rod moving along a prescribed curve and (b) the rod sliding longitudinally without fixed lateral slipping. Non-standard boundary conditions arising from constraints enable solving the equations. The one-dimensional equation captures system kinematics and dynamics without restrictive assumptions on trajectory or actuation, emphasizing the role of spontaneous curvature in driving and steering locomotion.
4.1.7 Bend Miscellaneous
Rods might exhibit interesting and mutable behaviors. Armanini et al. (2017) studied a soft robotic arm clamped at its base with a revolute joint, discovering that when the load is smaller than the buckling critical load, the rod behaves as an elastica compass. The closed curve can be computed analytically by the EBRT. Conversely, when the load exceeds the critical buckling value, the rod behaves like an elastica catapult, causing the snap phenomenon, where the rod releases very quickly elastic potential energy. They computed the dynamic motion using a simplified massless rod model and FEM and validated the behaviors by experimenting with an elastic rod.
Instead, Cacace et al. (2020) present a model for a simulated octopus tentacle based on the dynamics of an inextensible string with curvature constraints and curvature controls. They derive the EoM and appropriate boundary conditions characterizing the corresponding equilibria. The model gives fourth-order evolutive nonlinear controlled PDE, generalizing the classic Euler’s dynamic Elastica equation. They explore the reachability of the model and introduce and compare two optimal control algorithms based on the augmented Lagrangian method solved with adjoint-based gradient descent.
Subclass | References | Rod Model | Contribution |
Cable | Jones et al. (2009) | CRT | Static model for real-time shape estimation of CSM. |
Dehghani and Moosavian (2014b) | CRT | Dynamic model of a continuum finger for planar grasping. | |
Gao et al. (2016) | CRT | Model of notched continuum manipulator for MIS. | |
Pneumatic | De Payrebrune and O’Reilly (2017) | EBRT | Derivation of constitutive law of a planar bending actuator. |
Sadati et al. (2019) | EBRT | ROM and rod models with absolute and relative states validated on STIFF-FLOP. | |
Smart | Daily-Diamond et al. (2017) | EBRT | Dynamic model inspired by caterpillar motion to analyse ondulation mechanics. |
Pawlowski and Zhao (2018) | CRT | Modeling the coupling between CSM body and thermal actuation. | |
Parallel | Li et al. (2018) | EBRT | Shape modeling of a PCM made of soft panels. |
Tubular | Pattanshetti and Ryu (2019) | CRT | Estimation of stiffness variation on a monolithic tube for MIS. |
4.2 All Deformations
The second most popular deformation class concerns models accounting for all the primary deformations of continuum and soft robots. Below, we review the literature on rod models exhibiting all strains in cable-driven, pneumatic, parallel continuum, and biomechanical structures. Table 2 summarizes the class with the notable works.
4.2.1 Deformations by Cables
Renda et al. (2012) present a geometrically exact static Cosserat model of a conical tendon-driven CSM inspired by the octopus. Cable tension is modeled as a pure tension element, neglecting friction between the cable and the silicone body. After a FEM validation of the algebraic equations, the Cosserat model is validated experimentally for a static planar motion. The following work investigated the dynamic interaction of the silicone arm with a dense medium and the coupled tendon condition, addressing the external forces such as gravity, buoyancy, drag, added mass, and cable load (Renda et al., 2014). The dynamic experimental validation concerned the replication of characteristic octopus movements. In addition, an ablation study suggests that minimal internal viscosity guarantees solution stability and minimizes computational costs. Then, Renda et al. (2018b) introduced a unified dynamic Cosserat rod framework for underwater soft robots, which considers the nonlinearities of soft bodies, offering generality that extends to shells. The framework, which could aid the evaluation of potential soft robot designs, was validated experimentally using an autonomous vehicle with four cable-driven octopus-inspired CSM. At the same time, Renda et al. (2018a) presented an approach called piecewise constant strain (PCS) that discretizes continuous Cosserat rods in pieces, assuming constant strain for each piece. The PCS generalizes the PCC (Webster III and Jones, 2010) providing (i) numerical robustness given by the exponential formulation, (ii) inclusion of shear and torsion due to the CRT, and (iii) possible extension to rigid kinematic chains. The PCS is validated in real-world experiments with a cable-driven siliconic CSM.
Concerning discretization approaches, it was proposed a real-time discretization technique to solve the PDE of a Cosserat rod (Till et al., 2019). The method discretizes the time derivatives and solves the ODEs for the boundary value problem (BVP) in at each timestep. The authors validate the efficient method by comparing the simulation to high-speed footage of a cantilever rod.
For MIS applications, (Ibister et al., 2021) presents a dynamic CRT model for a continuum robot composed of a backbone and disks to support actuating tendons. The model includes the gravity (7) and tip wrenches produced by the tendons. The comparison between simulation and experiments shows a mean tip position error.
Recently, a cable-driven continuum robotic arm was integrated within an aerial manipulation system (Samadikhoshkho et al., 2020). The arm dynamics are modeled with a Cosserat rod with a floating base , incorporating both the actuation and quadrotor effects. They proposed a robust adaptive position controller and proved its stability using the direct Lyapunov stability theorem. The scheme is validated in simulation for quasi-static, passive, and active arm conditions.
4.2.2 Deformations by Pneumatic Actuators
Trivedi et al. (2007, 2008b) present the first dynamic Cosserat rod model for pneumatic CSM based on a fiber-reinforced model of the air muscle actuators. The model includes the effect of material nonlinearities, distributed forces, and payloads and is geometrically exact for large curvature, shear, torsion, and extension. The model is experimentally validated statically on the OctArm V for planar movements with vertical and horizontal base orientation, achieving results about ten times more accurate than the constant curvature model. Subsequently, the model is used within a model-based design optimization study for pneumatic and hydraulic CSM (Trivedi et al., 2008a). Recently, the PCS discretization was combined with Screw Theory, providing a new model of actuation sources for pneumatic actuators (Renda et al., 2017). This approach enables the routing of fluidic actuators (even tendons) along any path in the soft arm. They define a soft geometric Jacobian for the PCS formulation and a technique to identify the active DoFs, applying it to the study of manipulable and force-closure conditions. Finally, they validate the theory on a STIFF-FLOP arm (Cianchetti et al., 2013). Additionally, Roshanfar et al. (2021) described a pneumatic soft robot with a Cosserat rod model, which was solved as an initial value problem with homogeneous Neumann and Dirichlet boundary conditions for a given tip force. The experimental validation measured a tip position error of .
Analytical solutions and detailed FEM simulations are useful means for comparison and validation. For example, analytic solutions for the statics of a cantilever quasi-static Cosserat rod under external loads are derived (Grazioso et al., 2018). The position and orientation fields of the arm are defined on a Lie group, while the velocity and deformation fields are obtained from the time and space derivatives, respectively. A constitutive law between internal forces and strains enables the definition of the arm’s internal strain energy. Therefore, the principle of virtual work provides the static equilibrium equations. Finally, quasi-static rod configurations are compared to analytical solutions for pure bending and shear loads. Conversely, a comparative study for a Cosserat model and a FEM model of a soft pneumatic actuator was presented (Berthold et al., 2021), providing a preliminary parameter identification process. The FEM model was also validated experimentally. Moreover, an operational point-dependent static Cosserat model is obtained from a detailed volume FEM of a fiber-reinforced pneumatic silicone soft robot (Wiese et al., 2022). FEM simulations determine the actuation-dependent curvature, while the rod stiffness properties are identified by applying external forces to the FEM. The validation against FEM and experimental data, under various conditions with and without loads, show that characterizing a single module is sufficient to predict the behavior of a two-module version of the soft robot.
To tackle large deformations of pneumatic soft robots, Mishra et al. (2022) suggest using an explicit fractional-order Bouc-Wen model that considers material nonlinearities, such as hysteresis. The constitutive law incorporated in a Cosserat rod and tested on a Robotino-XT demonstrated superior accuracy to other models (Wang et al., 2022a).
4.2.3 Deformations with Parallel Manipulators
Till and Rucker (2017) addresses the problem of elastic stability in PCM, introducing a condition to evaluate the elastic stability of PCM solutions based on Kirchhoff and Cosserat rods. A numerical test evaluates this condition and a heuristic stability metric. After verifying that the stability test matches other classical results, they experimentally validate the approach on a six-link PCM. The same authors propose a generalized kinetostatics Cosserat rod model (Black et al., 2017). Model linearization provides the manipulator Jacobian, end-effector compliance, input stiffness, and wrench reflectivity matrices. This enabled the analysis of the PCM manipulability and force-sensing accuracy. Using a similar PCM, they achieve accurate actuation-based force sensing and low orientation error.
Conversely, Armanini et al. (2021) extended the PCS parametrization to soft PCM. They present a mathematical tool for modeling closed-chain geometries composed of soft and rigid elements. The formulation is applied to modeling and analyzing the Fin-Ray finger, especially for studying bending in the opposite direction of the external force. Furthermore, this formulation allowed the development of a novel design of the Fin-Ray finger, finding the optimal number of rigid ribs and their optimal orientation in terms of maximum payload and bending capability.
The elegant work of Lilge and Burgner-Kahrs (2022) present a CRT-based kinetostatic modeling framework for tendon-driven PCM, which enabled to efficiently solve the forward, inverse, and velocity kinetostatic problems. In addition, it facilitates the study of kinematic properties like reachability, singularities, manipulability, and compliance. Experiments on a PCM made of four tendon-driven backbones attached via spherical joints to a platform achieved median pose accuracy of and .
4.2.4 Deformations in Bioinspired Modeling
Cosserat rods can be useful tools for biomechanical studies. With application to flapping flight and passive swimming, Boyer et al. (2017) provide a general unified framework for modeling bio-inspired locomotion robots using soft organs. They model soft parts in two ways: (i) the floating frame approach and (ii) the geometrically exact approach. They implemented the two models in a unified general algorithm based on recursive Newton-Euler. The comparison of the two approaches revealed that the former needs fewer coordinates and stronger deformation approximations. Conversely, the latter is more accurate and manages better the external load. The validation of the unified framework simulated two bio-inspired examples of soft locomotion.
Recently, Zhang et al. (2019) realistically simulates musculoskeletal architectures in PyElastica. They assembled heterogeneous, active, and passive Cosserat rods into dynamic structures that mimic bones, tendons, ligaments, fibers, and muscle connectivity. The simulations involve biological and soft robotic case studies, including the human elbow joint, the design and fabrication of swimming and walking millimeter-long bio-hybrid bots, the synthesis of slithering snakes with intricate musculature layouts, and the replication of feathered bird wings - using up to rods. In these applications, the CRT demonstrated versatility to advance forward and inverse bioengineering designs and fundamental discoveries in the functioning of living organisms.
Boyer et al. (2022) investigates the relationship between CRT and optimal control theory, deriving the statics and dynamics of a continuum robot as solutions to a minimization problem. With this different formulation, the authors revealed significant theoretical insights, such as the intrinsic singularity of the dynamic simulation of continuum robots. Furthermore, the iterative solution of the optimal control problem corresponds to the continuous version of the Newton-Euler algorithm for rigid kinematic chains. Finally, the presented approach is extended to locomotors and applied to a bio-inspired continuum swimmer, validating the theoretical work through simulations.
Subclass | References | Rod Model | Contribution |
Cable | Renda et al. (2014) | CRT | Studied the dynamic interaction of an octopus-inspired CSM with a dense medium. |
Pneumatic | Trivedi et al. (2008b) | CRT | Model validation of CSM with fiber-reinforced actuators varying base orientation. |
Parallel | Till and Rucker (2017) | CRT | Studied and experimentally validated the elastic stability of PCM. |
Armanini et al. (2021) | CRT | Modeled closed-chain soft PCM with validation on a soft-rigid finger. | |
Lilge and Burgner-Kahrs (2022) | CRT | Studied reachability, singularities, manipulability, and compliance of PCM. | |
Bioinspired | Zhang et al. (2019) | CRT | Simulates musculoskeletal systems assembling heterogeneous active/passive rods. |
Boyer et al. (2022) | CRT | Studied the link between CRT and optimal control simulating locomotors and swimmers. |
4.3 Bend & Twist
Like bending, various robot types exhibit the combination of bending and twisting induced by diverse actuation types (Fig. 5). Table 3 summarizes the Bend & Twist deformation class.
4.3.1 Bend & Twist by Cables
Bending and twisting are usually induced by cables and tendons through specific routing paths (Fig. 5). Rucker and Webster III (2011) present Cosserat rod models for the forward kinematics, statics, and dynamics of a tendon-driven continuum manipulator. The curved tendon paths deform the body in various shapes. Starting from the classical CRT, they derive the expression of distributed loads applied by tendons on the backbone . The model validation considers straight and non-straight tendon routings, point loads, and distributed external loads . The simulated matched experimental tip positions with a 1.7%L average error. Another experimental work presented a continuum arm with a Nitinol backbone and three tendons turning around the centerline with a constant radial distance (Dehghani and Moosavian, 2013), which extended a previous setup lacking an actuation system (Dehghani and Moosavian, 2011). They modified differential terms of an existing CRT (Rucker and Webster III, 2011), retaining the original accuracy with reduced numerical effort. The model was extended to real-time control using a stable moment-based algorithm to solve the BVP considering the physical characteristics of the rod (Dehghani and Moosavian, 2014a). They also propose a method to reduce the cumulative numerical integration errors that could make the rotation matrices non-orthogonal. Finally, they use the model to develop a Jacobian-based quasi-static tip position controller for a multi-segment continuum arm.
Cosserat rods can simulate the mechanics of several robotic systems, including robotic tails and octopus-inspired arms. The work in (Rone and Ben-Tzvi, 2014) used a Cosserat rod to study a cable-driven multi-section continuum robotic tail envisioned for onboard mobile robots. Prescribed tail configurations or trajectories are analyzed, while the governing equations compute the loading at the tail base transmitted to the mobile robot. This study investigates the impact of trajectory and design factors on the resulting loading profiles, considering factors like mode shape, speed, bending magnitude, and bending plane angle. Key findings include that a non-straight tail reference configuration increases the moment magnitude, short proximal segments provide a larger tip motion for a given bending angle, and higher mass distribution towards the final nodes enhances bending moments and control fidelity. Xu et al. (2017) propose a Cosserat rod model to describe the bending and torsion behavior of a bionic tendon-driven octopus tentacle. Given the location of the four cables, the model maps cable tension to expected bending. Exploiting the derived mapping, they propose an open-loop controller for the robot. Similarly, Niu et al. (2019) derive closed-form equations of a soft robot modeled through the CRT. The equations are validated with a cable-driven soft arm inspired by an octopus tentacle, reaching accurate predictions of the end-effector positions. Chikhaoui et al. (2019) investigated the kinematic modeling of a magnetically extensible tendon actuated continuum robot (METAbot) (Nguyen and Burgner-Kahrs, 2015). The comparison between a PCC beam mechanics approach and a Cosserat rod shows that the latter has higher position accuracy, although computationally heavier. Recently, Janabi-Sharifi et al. (2021) presented a tutorial on the dynamic Cosserat rod modeling of tendon-driven continuum robots. The tutorial explains the core assemblies of tendon-driven continuum robots. Then, it provides a unified approach for the dynamic analysis of tendon-driven continuum robots that bend and twist, including multi-section, concentric tubes, and parallel cooperative robots.
4.3.2 Bend & Twist in Soft Pneumatic Actuators
Actuation-induced torsion is not common in fluidic actuators. However, a Cosserat rod model for a CSM with two asymmetric pneumatic actuators that bend and rotate was introduced (Uppalapati et al., 2018). They optimize the rigidity parameters and the curvatures for quasi-static shapes. The model reproduced the experimental static workspace with a tip load attached. The robot’s dexterity was enhanced integrating an additional soft rotating actuator (Uppalapati and krishnan, 2021). The CRT predicts the deformed robot configuration subjected to external loads, and they experimentally determine elasticity and actuation parameters.
Another study introduced a multi-segment Euler-Bernoulli rod model to predict the bending deformation of micro-tube pneumatic actuators (Ji et al., 2019). They first employ the line-segment model and then extend it to a multi-segment version of Euler-Bernoulli’s rod model for a rapid design, analysis, and optimization of the bending deformation of such actuators. Finally, an analytical model of a generalized PneuNet soft actuator combines the minimum potential energy method and an Euler-Bernoulli rod (Gu et al., 2021). The study compared the rod model to FEM simulations and experimental results, achieving accurate deformed shape predictions, given the geometry and the applied pressures.
4.3.3 Bend & Twist in Tubular Robots
Biomedical applications could leverage CTR and eccentric tube robots that bend and twist. A geometrically exact KRT solved the statics problem for CTR, describing the shape of a collection of concentric pre-curved tubes under general wrench distributions (Rucker et al., 2010a). The model is validated experimentally for an active cannula with three concentric, telescopic, pre-curved Nitinol tubes in several configurations with varying tip loads. The results establish the basis for model-based position and force control of active cannulas and their use as force sensors under external loads. Also, they tackle the integration of unmodelled effects such as friction, elongation, and transverse shear (Rucker et al., 2010b). Subsequently, Till et al. (2020) present the first dynamic model for CTR. The model describes the coupled inertial dynamics of pre-curved concentric tubes considering the dynamic effects of linear and rotational tube inertia, material damping, Coulomb and viscous friction, and the inertia of a rigid body at the robot tip. The experimental validation measures the dynamic behavior during tissue grasping and snapping.
The structure of the ETR extends the CTR design. A kinematic model of a two-channel ETR with two continuum arms delivered through the channels is presented (Wang et al., 2019). Pre-curved superelastic tubes lining the arm’s working channels produce sheath deflection. Cosseat rods model the pre-curved tubes and the central backbone, achieving promising tip pose predictions for pure rotation actuation. Similarly, a continuum ETR sheath for MIS such as neuroendoscopy, requiring delivery of multiple robotic arms, was presented (Wang et al., 2021a). Super-elastic tubes lining the arm channels actuate the sheath, while distributed disks are attached to the backbone. The tube kinematics and the robot backbone are modeled using the CRT. Three experiments with a two-arm sheath validate the tip position accuracy. Another work proposes a kinematic Cosserat rod model of a snake-like continuum robot for MIS (Wang et al., 2020). The robot geometry comprises three identical super-elastic Nitinol tubes held eccentrically by a surrounding sheath. Assuming sheat inextensibility, they show that different actuations virtually yield the same robot shape and tip pose. Simulations and experiments validate the model for a single-stage ETR.
Safety is crucial in surgical procedures. A method to compute compliance of continuum robots, modeled as quasi-static Cosserat rods under diverse loads, was devised (Smoljkic et al., 2014). The robot compliance is described by ODEs and integrated with the forward kinematics. The experiments involve a non-actuated flexible rod composed of a super-elastic Nitinol tube surrounding a braided Pebax tube. The rod mounted on a force sensor with position sensors embedded in the tip is subject to external tip forces. A least square fitting of the measured tip trajectory versus the direct kinematics identifies material parameters and tube configuration. Finally, they compared the velocities obtained through position differentiation against those obtained multiplying the differentiated forces and the computed compliance matrices.
4.3.4 Bend & Twist with Smart Actuators
The KRT modeled TCP emphasizing critical factors like coil load and twist and incorporating them into the actuation equations for TCP muscles (Wu and Zheng, 2020). The model captures the quantitative relationship between tensile actuation and fabrication load, serving as a predictor for tensile stroke. Also, it was validated through TCP muscles made of silver-coated Nylon fiber under different fabrication loads, measuring recovery stress and maximum stroke during heating. Thermo-mechanical tension tests conducted at various room temperatures identified the elastic modulus of the fibers. To achieve variable stiffness in continuum robots, Wang et al. (2022b) propose to use winding-styled ropes inspired by snakes’ behavior. Controlling the temperature of an SMA spring increased their stiffness up to . The experimental validation of a kinetic-static Cosserat rod model obtained an error below .
4.3.5 Bend & Twist in Parallel Manipulators
4.3.6 Bend & Twist with Buckling Instability
Buckling instability is an important factor to consider in soft robots. With applications in soft robotics, Sipos and Várkonyi (2020) explores the interplay between spatial buckling and arbitrary intrinsic curvature on a simulated clamped Cosserat rod with a vertical tip load. The aim is to tune the control parameters to find the highest load for a given free-end position or the maximum reaching of the cantilever in a direction for a given load. The numerical investigation of the spatial instability of in-plane configurations of soft elastic cantilevers subject to an endpoint load demonstrates the importance of out-of-plane instability. Moreover, the capability of controlled intrinsic curvatures to stabilize the rod in configurations of high reaching is shown. Significant improvements occur for flattish cross-sections where the rigidities associated with out-of-plane bending and torsion are larger than the in-plane stiffness. At the same time, arms with high and narrow cross-sections are slightly more performant. Buckling instability can be triggered. Forghani et al. (2021) analyze the inverse dynamics and control of a bacteria-inspired uniflagellar robot in a fluid medium at a low Reynolds number. The robot, mimicking the locomotion of flagellated bacteria, features a flexible helical filament (flagellum) connected to a spherical head. A Kirchhoff elastic rod describes the flagellum, which rotates about the head at a controlled angular velocity and generates propulsive forces for the forward motion. The results demonstrate that by triggering the buckling instability of the flagellum, the robot follows a prescribed path in space. The control scheme only involves the flagellum angular velocity, and a deep neural network captures the dynamics predicting the robot’s trajectory. However, higher maneuverability tasks like drug delivery require multiple flagella, advanced controllers, and control inputs.
4.3.7 Bend & Twist in Insect-like Wings
The applicability of rod theories goes beyond manipulators. A study presented a hybrid dynamic Cosserat rod model of a 3D aerial insect-like robot with two soft wings (Belkhiri et al., 2012). In particular, Cosserat rods described the leading edge of the wings, which bend and twist based on continuous Newton-Euler dynamics. The wings, connected to the insect thorax through revolute joints, are treated as rigid bodies. The hybrid model combines discrete and continuous models considering the curvature deformations and external aerodynamic forces. The model replicates the kinematics and internal deformations of the wings and the rigid net insect motion produced by a flapping flight.
Subclass | References | Rod Model | Contribution |
Cable | Rucker and Webster III (2011) | CRT | General dynamic model of a continuum manipulator with curvilinear actuators. |
Pneumatic | Uppalapati and krishnan (2021) | CRT | Design and modeling of a CSM with parallel asymmetric fiber-reinforced elastomers. |
Tubular | Till et al. (2020) | CRT | Dynamic model of CTR validated for tissue grasping and snapping. |
Wang et al. (2021a) | CRT | Modeling of a superelastic ETR sheath for biomedical procedures. | |
Smart | Wu and Zheng (2020) | KRT | TCP muscles incorporating coil load and twist into actuation model. |
Buckling | Sipos and Várkonyi (2020) | CRT | Study relationship between buckling and on simulated cantilever rods. |
4.4 Bend & Stretch
The simultaneous bending and stretching emerge in many cable-driven and pneumatic CSM (Fig. 5). Table 4 reports illustrative examples of the Bend & Stretch class.
4.4.1 Bend & Stretch by Cables
Deflection and stiffness estimation are crucial for continuum robots actuated by cables. For instance, robot deflection and stiffness were predicted considering environmental loads with constant axial tendon displacements (Oliver-Butler et al., 2019). Preliminary experiments with an Euler–Bernoulli rod evaluate how tendon stretch and routing path affect the robot stiffness at any point. The analytical results enable stiffness analysis of candidate robot designs with light simulations and suggest the ability to increase robot stiffness by using converging tendon paths. Finally, they extend the study to a non-linear Cosserat rod model with large deflections, handling prescribed tendon displacements, tendon stretch, pretension, and slack. Experimental validation on a prototype shows good agreement. Also, a novel stiffness modulation approach using a Cosserat rod model was proposed for a tendon-driven soft CTR (Xiao et al., 2022). The experimental validation shows an accurate shape estimation. They also demonstrate that the robot stiffness changes with a Nitinol backbone. In particular, the tip compliance can be halved, with potential applications in manipulation tasks.
Concerning workspace estimation, the CRT is also applied to investigate equilibrium positions reachable by the end-effector of a planar tendon-driven CSM with a discrete actuation space (Walid et al., 2022). Experiments on a physical prototype efficiently estimated the robot workspace. Another example predicts the soft arm shape combining the EBRT, which considers the influence of external forces in bending, with the PCC model, which efficiently derives a relationship between cable length and arm curvature (Chen et al., 2022). Experiments on a cable-driven soft arm subject to external forces, or , yield a maximum error in the range .
4.4.2 Bend & Stretch in Soft Pneumatic Actuators
A study compares five models, including the Cosserat rod, a dynamic lumped system, and approaches with constant or variable curvature (Sadati et al., 2017b). The comparison made in simulation and on the STIFF-FLOP soft arm (Cianchetti et al., 2013) demonstrated that the lumped mass model better predicts the dynamics, while constant curvature methods are less accurate but are suitable for design optimization. Additionally, they modeled CSM actuated by braided pneumatic actuators (Sadati et al., 2017a). This specific design causes cross-sectional deformation, which was captured analytically through geometry deformation and CRT. This approach outperformed constant curvature approaches by and variable curvature by . A subsequent work presents a semi-analytical model for the dynamics of multi-segment CSM (Sadati et al., 2017c). The Cosserat rod static and Lagrangian dynamic models are derived and combined with the Ritz-Galerking method to reduce the infinite state space of CSM to a small finite number of states.
Moreover, a static Kirchhoff rod model was proposed for a hybrid CSM characterized by pneumatic actuators with embedded elastic rods, computing its deformations using the minimum energy principle (Sun et al., 2019). In addition, the classical CRT modeled a variable-length pneumatic CSM (Gilbert and Godage, 2019). They optimize the cross-sectional area of the actuators and a parameter relating stiffness and damping in the constitutive law. The model is validated experimentally on dynamic bending deformations obtained by pressurizing the actuators with alternating square pulses. A study characterized the bending behavior of a cantilevered extensile fluidic artificial muscle (Garbulinski and Wereley, 2022). An Euler-Bernoulli rod determined static properties like the bending stiffness, which increases more than four times in the pressure interval psi. Experimental data are applied to assess the dynamic properties, such as natural frequency and damping. Also, it was developed an analytical model of a PneuNet actuator based on the Euler-Bernoulli finite strain rod (Sachin et al., 2022). They describe the actuator behavior in free motion and for tip contacts and accurately predict the actuator deformation and the force characteristics. Another work showcases a parallel extensible soft robot with three pneumatic actuators, each with a silicone body, inner air chamber, winding wiring, and sealing head (Wang et al., 2022c). The dynamic model of a single actuator uses the CRT, and an interactive constraint topological structure solves the constraint force and moment of the parallel robot. A linear mapping between the pressure and the extension force is established for each actuator. Validation measures the static step response of a single actuator and conducts dynamic experiments with the complete soft robot. The model accurately captures the actuator dynamics and the parallel robot constraints. Rod models met traditional robotics models where the EBRT augmented the kinematic model of a pneumatic soft bending finger treated as a planar serial rigid robot parameterized with the Denavit-Hartenberg convention (Flores-Martínez et al., 2022). The experimental validation compares the actuator trajectory versus the kinematic model and the flexion-extension motion trajectory of average human fingers.
Manufacturing uncertainty could be significant in fluidic soft actuators. For instance, Eugster et al. (2022) described the kinematics of a soft arm using a CRT with a nonlinear pressure-dependent constitutive law, using the principle of virtual work and modeled the actuator with strain energy functions. Manufacturing imperfections are considered by scaling the relations of extensional and bending stiffness and by a constant shift of the actuator position . Assuming the pressure dilates the chambers, the model also considers cross-section deformations with a pressure-dependent diameter . The model is validated experimentally for static stretching and bending motions in vertical and horizontal mountings. They compare models with constant or pressure-dependent chamber radius, with or without chamber repositioning on the cross-section, and with a linear or Ogden material law. The strain energy for an Ogden material (Odgen et al., 2013) is
Here, is a material parameter, and is the principal axial stretch. Similarly, Alessi et al. (2023a) presented a dynamic Cosserat rod model for a 3D-printed pneumatic CSM. The model considers the pressure transient with a parameterized smooth step. Also, it captures the manufacturing variability of the actuators using strain gains that tune the pressure-strain relation for each actuator
where is the interior actuator area. An extensive experimental validation includes quasi-static and dynamic motion patterns at various actuation frequencies. Finally, an ablation study evaluates the contribution of different model components, uncovering that neglecting gravity and manufacturing uncertainties could degrade performance up to , depending on the motion.
Concerning fiber-reinforced actuators, Berthold et al. (2022) introduces a Cosserat rod model to deepen the understanding of the pressure-induced deformation behavior of a fiber-reinforced soft pneumatic actuator with three DoFs. By converting the inner chamber pressure into an equivalent force, they consider the compression effects of lateral pneumatic chamber walls. The validation compared the results to FEM simulations, obtaining a maximum mean tip position error of . Differently, Hanza and Ghafarirad (2023) introduces an inhomogeneous Cosserat rod model for a CSM with pneumatic fiber-reinforced actuators, using a nonlinear strain-force relationship to account for the effect of radial pressure. The proposed method outperformed the standard CRT by about .
Alongside, Liu et al. (2023) discusses another modelling approach of a pneumatic fiber-reinforced soft actuator with a semicircular cross section. Here, the EBRT is employed to explore the effects of self-gravity and external loads on the actuator configuration, to predict its shape.
Dou et al. (2023) introduce a hybrid CSM with pneumatic chambers and an antagonistic rod-driven compliant mechanism enabling elongation, contraction, bending, and stiffness regulation. They derive a kinematic model by adapting the EBRT and demonstrate the benefit of the compliant mechanism with an experimental validation comprising way-point tracking.
Recently, Lamping and de Payrebrune (2023) presented a study on soft robotic actuators that utilize pneumatic actuation with multiple chambers. The modeling approach uses the EBRT to characterize the deformation, providing a practical method for accurately determining the rod parameters essential for predicting the performance of these actuators, enhancing their design and control.
Subclass | References | Rod Model | Contribution |
Cable | Oliver-Butler et al. (2019) | EBRT/CRT | Stiffness analysis for various tendon stretch and routing paths . |
Chen et al. (2022) | EBRT | Augments the PCC with a rod model to quickly solve the CSM deformation under . | |
Pneumatic | Sadati et al. (2017c) | CRT | Utilizes the Ritz-Galerkin method to reduce the continuous state space of rod models. |
Alessi et al. (2023a) | CRT | Ablation study of model addressing manufacturing uncertainties in a 3D-printed CSM. | |
Dou et al. (2023) | EBRT | Design and modeling of a hybrid CSM incorporating an antagonistic compliant mechanism. | |
Smart | Goldberg et al. (2019) | CRT | Simulated locomotion of a caterpillar-inspired soft robot actuated by SMA. |
4.4.3 Bend & Stretch with Smart Actuators
Regarding locomotion with smart actuators, Goldberg et al. (2019) presents a model for a caterpillar-inspired soft robot actuated by SMA, leveraging the work by (Bergou et al., 2008) related to the elastic rod theory and the Lagrange motion equation for systems of particles. A planar version of Bergou’s formulation is employed here to model the dynamics of a single SMA actuator. The actuator stiffness and damping parameters are first estimated and subsequently used to simulate the locomotion of the caterpillar robot fabricated from a series of SMA actuators. Interestingly, increasing the actuator stiffness by yields simulation displacements higher than the prototype by .
4.4.4 Bend & Stretch in Parallel Manipulators
The CRT also aided the design, statics modeling, and workspace analysis of an extensible rod-driven PCM with five DoFs at the tip (Wu and Shi, 2022a). A variant of the Levenberg-Marquardt method makes the PCM model suitable for real-time applications. Open-loop errors on the tip position in tracking tasks are lower than in PCC models. Also, they propose a numerical procedure for efficiently estimating partial workspace using a ROM.
4.5 Bend & Twist & Stretch
Five contributions explored bending, twisting, and stretching combined. For cable-driven actuation, it was presented a general geometrically exact static CRT for CSM (Renda and Laschi, 2012). The work highlights the role of torsion deformation. The variety of the design parameters allows the simulation of a wide range of continuum manipulators and the coupled tendon drive for multiple sections. Similarly, a cable-driven conic CSM made of silicone was modeled by combining a geometrically exact dynamic Cosserat rod and the Kelvin model of viscoelastic materials (Wang et al., 2016). An experiment with the arm fixed at the base and oriented horizontally, subject to gravity, without actuation (), validates the viscoelastic model, which matches the experimental data better than a pure elastic model. Finally, the validation executed 2D and 3D motions actuating different cable combinations (). Also, a discrete Cosserat rod combined with a FEM described the dynamics of cable-driven soft robots (Adagolodjo et al., 2021). In particular, they use FEM to model the robot structure and Cosserat for the robot actuation. The implicit backward Euler time stepper integrated the differential equations for numerical stability. Direct and inverse simulations demonstrated the suitability of the approach for modeling robots actuated by cables or rods.
Aimed at colonoscopy, the CRT was employed to reconstruct the shape of a continuum endoscope given the force exerted on the distal end (Ryu et al., 2018). The experimental validation shows that longer endoscopes produce higher tip errors.
Also, Cosserat rods modeled a six-link PCM (Ghafoori and Khalaji, 2020). The model, initially validated for a single rod obtaining an average tip position error of , is applied to a six-link robot with good performance.
In a recent work (Roshanfar et al., 2023), the authors modelled and experimentally validated a soft robot with combined pneumatic and cable actuation. The robot is modelled using CRT where the tangent elastic modulus is expressed as a function of the internal pressure. In addition, the hyperelastic constitutive model was integrated into the Cosserat framework.
4.6 Bend & Stretch & Shear
Two works combine bending, stretching and shearing. For multi-backbone continuum robots, it was proposed a variable curvature statics-kinematic Cosserat rod model incorporating multi-backbone structural constraints (Chen et al., 2021). The model is validated on a robot with two multi-backbone continuum segments, each with four nitinol rods. The push-pull actuation bends the continuum robot. The experiments evaluate the tip positioning accuracy for circular trajectories and the shape discrepancy when the robot is subject to an end-point load in two configurations. They also implement the inverse kinematics in a real-time open-loop controller for a rectangular path following, with and without tip load. The proposed model outperforms the PCC.
Differently, a hybrid model combining the classical CRT with a data-driven stiffness estimation was developed (Bartholdt et al., 2021). Rather than solving numerically nonlinear constitutive equations, the model relies on a work point-dependent linear stress-strain relationship for computational efficiency. Experimental validation is carried out by identifying the stretching and bending stiffness of a soft pneumatic actuator and learning the actuation to stiffness maps using artificial neural networks.
4.7 Bend & Twist & Shear
The only work that performed an experimental validation for the compound class of bending, torsion, and shearing is by Grazioso et al. (2019), who proposes a novel FEM-inspired spatial discretization technique. In particular, the rod is divided into several nodes, defining a helicoidal shape function for the interpolation. Thanks to the specific choice of the shape function, the Cosserat rod dynamics was written with Lie Groups. The method, implemented in a new simulator called SimSOFT, is validated on the pure bending of a cantilever soft arm and the pure in-plane rotation of a soft arm with varying external conditions. In addition, they performed an experimental validation on the Princeton arm to benchmark the coupling of bending, torsion, and shear. The validation results report excellent performance in terms of accuracy and computational efficiency.
Bentley et al. (2023) employ a Cosserat string formulation to derive a tissue and needle force model. It describes normal and frictional forces along the shaft as a function of the planned needle path, friction model and parameters, and piercing force. The proposed force model and associated cost metric are safer and more clinically relevant for motion planning. They fit and validate the model through physical needle robot experiments in a gel phantom. The force model defines a bottleneck cost function for motion planning, which is evaluated against a path-length cost function in random environments, achieving a 62% peak force reduction.
4.8 Stretch & Shear
Combined stretching and shearing deformations are characteristic of peristaltic locomotion (Fig. 5). The peristaltic locomotion of earthworms was studied with a continuous model of compressible and incompressible slender bodies represented as Cosserat rods (Hemingway and O’Reilly, 2021). Incompressibility is enforced as an internal constraint using Green and Naghdi’s theory of a directed rod. Two linearly elastic isotropic material models are assumed. The material parameters are identified from experiments for small deformations. Motion effects from actuators or muscle contraction are modeled as external compressive loads using a doublet function for an assigned centerline force and a uniform pressure for a pair of assigned director forces. The method is showcased in a simulation of a soft robotic device.
4.9 Stretch
The stretching deformation successfully simulated peristaltic locomotion. The forward motion results from wave propagation through contraction and extension of the robot segments along the movement direction (Fig. 5). With bioinspired robotics applications, Nguewou-Hyousse et al. (2019) modeled a caterpillar soft robot using a network of linear and torsional springs and massless rods, modeled as planar discrete elastic rods (Goldberg et al., 2019).
4.10 Discussion on rod models
The analyzed works demonstrate a steady interest in rod models for continuum and soft robots. As shown in Fig. 6, the number of published papers on rod-based models increases.
4.10.1 On the Deformations
Researchers developed numerous rod-based models reformulating and adapting classical rod theories in the soft robotic field. Herein, we reviewed diverse models, which described different combinations of strain modes, giving rise to nine deformation classes (Fig. 4).
The most common class in literature is Bend, with the of the examined papers. This popularity is due to the propensity of slender bodies to bend. In fact, due to the assumption of , with cross-section diameter, the bending stiffness turns out smaller than the stretching or shear ones, easing the bending deformation. In addition, most used actuator paths are linear, which excite predominantly bending modes. These reasons justify the approximation of considering only the bending mode. Furthermore, we can observe that most models in the Bend class leverage the EBRT. However, as summarized in Table 1, several models employed the general CRT. While this represents a computational surplus for trivial bending structures, the generalization of Cosserat models pays off if the model gets deployed in physical interaction tasks.
The next class is All Deformation, with the of the examined papers. As expected, only models based on the CRT achieve all deformations (Table 2). This group of models received a deep study in the last decade, thanks to the rising interest in physical interaction with unstructured environments where all strain modes are significant. Indeed, these models are useful to describe accurately the deformations of slender robots caused by friction (e.g., anisotropic friction for locomotors) or by different mediums (e.g., water). In addition, the most common actuation sources (i.e., pneumatic and cables) excite elongation and compression, causing a significant change in the dimension workspace.
The third class accounting for of the investigated studies is Bend & Twist. These models mainly describe robots in which the unshearability and the unstretchability approximations hold, as summarized in Table 3. Including the twisting deformation allows the representation of backbones with generic 3D curves. However, the twisting deformation can be excited only by specific actuator paths (e.g., helicoidal cables) or asymmetric actuators.
Another significant class, summarized in Table 4, is Bend & Stretch () for the wide diffusion of linear extensible soft pneumatic actuators and tendon-driven robots. Such actuators extend or compress upon solicitation, which induces the continuum body to stretch and bend thanks to their radial disposition on the cross-section.
The remaining five categories are less explored. While the Bend, Twist & Stretch () is a particular case of the All Deformations class, the other classes comprehend few works. The lack of work is due to the difficulty of exciting these specific strain modes with the current actuation technologies or the small number of applications in which the assumptions are valid. In addition, fewer studies that explored shearing per se could be due to the tendency to consider shearing an undesirable or negligible deformation. However, shearing could be significant when there is a prominent environment interaction, as observed in peristaltic locomotion (Hemingway and O’Reilly, 2021) or complex manipulation (Grazioso et al., 2019).
4.10.2 On the Spatial Discretization
One characterizing feature of rod models is the spatial discretization method, which significantly impacts the model’s accuracy and complexity. For simulation and control purposes, the discretized model must be computationally efficient and as accurate as possible. In the analyzed literature, exist two main discretization techniques: Strain Parameterization (Sec. 2.2.3) and Discrete Elastic Rod (DER) (Sec. 2.2.4). The former is based on a discretization directly in the configuration space, allowing the user to choose the degrees of accuracy. In addition, it is possible also to neglect specific strain modes to reduce the computational time. The latter discretizes the length of the rod, considering also the external forces applied to the nodes. Furthermore, most of the existing spatial discretization algorithms are already implemented in dedicated simulators, such as SoRoSim (Mathew et al., 2022) for SP and PyElastica (Naughton et al., 2021) for DER. Straddling geometric and mechanical models, some examined works proposed to fit discrete points with functions, such as PH or Euler’s spirals, which is useful, especially for Shape Estimation.
4.10.3 On the Identification and Validation
Identification and validation of rod models play a crucial role in accurately reproducing the experimental robot behaviors. Despite the high computational cost, it emerged that detailed FEM simulations served a dual purpose:
to assess the correctness of the rod model,
to estimate the unknown parameters.
In the absence of precise FEM models, lumped-mass models with a high number of DoFs were employed. Otherwise, researchers resort to model optimization to match experimental data. In particular, when the robot morphology is not uniform, researchers usually estimate geometrical properties (e.g., cross-sectional area ) and material characteristics (e.g., elastic modulus , damping coefficient ). Experiments typically conducted are with the robot at different base orientations subject to external loads or at the tip. Most rod models are identified and validated through motion data collected with predefined actuation profiles . However, the literature needs rigorous mechanical tests to yield more accurate models, especially for unstructured physical interactions with the environment. Concerning performance metrics, rod models predicted experimental tip positions with Euclidean errors in the range . However, only a few studies assessed the angular errors, which play a role in dexterous manipulators. In summary, the wide use of experimental data suggests an increasing need to reduce the reality gap to develop more accurate controllers.
4.10.4 On the Impact of Materials and Manufacturing
Most of the analyzed rod models employed continuum and soft robots built with diverse materials, including traditional elastomers (e.g., silicone, rubber) or other polymers (e.g., SMA, hydrogels). Each material offers distinct properties, influencing the robot’s compliance, flexibility, and functionality. However, effectively modeling the nonlinear behavior inherent in soft materials using rod theories presents challenges. For instance, some combinations of materials and actuation mechanisms cause notable changes in cross-section areas or volume, which may render the rod-like assumption overly restrictive. Nonetheless, an extension of the CRT considering geometrical rescaling (15) could approximate complex dynamics (16).
The viscoelastic constitutive law of the material impacts significantly on the accuracy. In the classic rod theories, this relation is supposed linear, as already shown in (6). However, especially in the case of pneumatic actuation, the linear constitutive law is no longer valid. To tackle this issue, many analyzed works presented models with a nonlinear constitutive law, using hyperelastic model such as Odgen.
Manufacturing techniques provided continuum and soft robots with unprecedented dexterity (Wallin et al., 2018). However, manufacturing uncertainties may impact model accuracy. Recent rod models addressed manufacturing uncertainties considering variations in material (Alessi et al., 2023a) and geometric properties (Eugster et al., 2022). Despite these advancements, some gaps persist, particularly in capturing hysteresis effects or integrating the dynamic restoring process of self-healing materials into rod-based models. Hysteresis, a prevalent phenomenon in soft materials, strongly impacts the accuracy and predictability of the robot’s behavior. Furthermore, incorporating self-healing materials can potentially enhance the durability and longevity of soft robots but remains relatively unexplored in rod theories (Terryn et al., 2017). Addressing these challenges could advance the capabilities of continuum and soft robots, enabling their integration into diverse applications and real-world tasks.
5 Controllers using rod models
Controlling continuum robots can be very challenging, due to the intrinsic features of this type of robots. Material nonlinearities (e.g., hysteresis), the infinite-dimensional configuration space, and the under-actuated nature of the system make the direct implementation of well-known controllers for rigid manipulators difficult or impossible.
The literature on the control of rigid robots primarily focuses on controlling the end-effector in terms of pose, velocity, or force. In the realm of continuum robots, the natural equivalent of this task is tip control. To exploit the continuum nature of soft mechanisms and achieve tasks that are impossible for their rigid counterparts, the researchers have proposed two other tasks: Multi-Point (MP) control and shape control. The goal of MP control is to control the pose of a set of cross-sections, while shape control aims to regulate the backbone of the robot, fully utilizing the potential of the continuum structures.
To tackle the inherent properties of continuum robots, rod theories have been widely used for control design, resulting in an increasing trend of works in the last decades, as shown in Fig. 6.
The developed controllers fall into two different categories: model-based and learning-based.
5.1 Model-based Controllers
The Model-Based (MB) control strategy involves leveraging prior knowledge of the robot to guide it toward achieving a specific task. The motivation behind developing MB controllers is the possibility of mathematically guaranteeing the stability and performance of the controlled system. However, implementing controllers in physical prototypes can be challenging, often requiring an extensive identification procedure. This context sees the spatial discretization technique as crucial to balance accuracy and computational efficiency. Therefore, the model selection and the discretization technique are integral parts of the control design process.
5.1.1 Formalization
To formalize the control problem, let us consider a generic configuration vector, whatever discretization algorithm is used. According to Della Santina et al. (2023), the shape control problem can be particularized in two sub-problems: (i) regulation and (ii) trajectory tracking (TT). In the last case, the goal of shape control is to find an input such that
where are the desired configuration and its time derivative, respectively. The condition (48) can be particularized in the regulation case, imposing .
Similarly to rigid robots, it is possible to formalize the tip control problem, rewriting (48) in the task space, such as
where and are the desired pose and velocity twist of the tip. Similarly, the tip pose regulation condition can be obtained, particularizing (49) with . Lastly, the MP control problem can be formalized by extending (49) in a set of cross-sections’ desired poses and velocity twists.
To solve these control problems, it is possible to represent the MB control framework in a general scheme, shown in Fig. 7. The block “Sensors” refers to the sensors mounted on the robotic platform, mapping in the measurements . The “Estimator” block implements the Shape Estimation/Reconstruction algorithms, providing an estimation of , , defined as , . Finally, the MB controller block computes the input , processing the error between the estimated configuration and the desired one. More in-depth, the control input is composed of two actions (Fig. 7). The former is the feedforward action , depending on the chosen model, the estimated state and the desired trajectory. The latter is the feedback action that depends on the error w.r.t. the desired configuration and its time derivative. Finally, the specific control algorithm computes the control output combining the feedforward and feedback terms.
5.1.2 Inverse Kinematics
One of the first approaches proposed in the MB literature is the Inverse Kinematics control. The main idea is to use the inverse kinematic model to find the actuation input such that the tip reaches a desired pose or velocity twist. In the above case, the rod-based kinematic models enable the design of controllers that account for the continuum robots’ deformations.
In the planar case, the EBRT provides a simple and efficient kinematic model. Ataka et al. (2020a) proposed a kinematic pose control for an inflatable eversion CSM for trajectory tracking in a planar environment. Assuming a pure bending robot, they used EBRT to model the effect of internal inflation pressure on each bending segment of the prototype. A further advancement in the modeling is the characterization of the variable structural stiffness caused by changes in inflation pressure. Additionally, an observer based on the Extended Kalman Filter estimates the unknown model parameters online from chamber pressure and bending angle measurements. These parameters approximate the actuator-space Jacobian matrix online, which controls the CSM’s tip pose under various stiffness conditions. Simulation and experimental validation using an inflatable CSM made of unstretchable fabric demonstrated that the proposed method achieved position control in a planar environment despite different robot inflation pressures (i.e., stiffness values).
Mauzé et al. (2021) achieve micrometer positioning accuracy with a PCM made of three planar kinematic chains linked to a rigid moving platform. Each chain has a prismatic actuator and a flexible rod that deforms continuously, with bending and twisting deformations modeled with the CRT. A regular Cartesian controller based on inverse kinematics is implemented using the parameters identified for the forward kinematic model. The controller is validated for a circular and square trajectory. For both, large orientation displacements of the platform increase the angular error.
A closed-loop multi-point orientation control strategy for discretely magnetized continuum manipulators was developed (Richter et al., 2021). A non-homogeneous magnetic field model conforming to Maxwell’s equations was formulated and applied to the actuation system to predict the exerted wrenches on the manipulator. The approach is demonstrated on a manipulator containing two permanent magnets actuated inside a non-homogeneous magnetic field. Here, a quasi-static CRT models the manipulator deformation under external wrenches, while a Jacobian provides the actuation inputs. Experiments in simulation and on a real robot demonstrate that the algorithm independently controlled the orientation of two interconnected magnets in a non-uniform magnetic field. However, the control frequency is limited to Hz due to the computational cost of shape reconstruction and the Jacobian.
Campisano et al. (2021) propose a controller for CSM under tip follower actuation. In particular, the strategy is to couple a quasi-static regime and sensor feedback of the actuation sources. This approach achieves (i) rapid computation of the Cosserat rod model, (ii) actuator nonlinearities compensation, and (iii) oscillation reduction. In conclusion, the closed-loop control algorithm is validated experimentally and compared with an open-loop one, to show the effectiveness of the design process.
Renda et al. (2022) proposed an inverse kinematics controller based on the GVS model, using the actuators-strain mapping in (39). The control algorithm is designed for trajectory tracking of a set of points in . The kinematic model is extended, including a map between the length of threadlike actuators and the configuration of the CSM. This formulation covers the general case of actuators’ routing, such as discontinuous or overlapping routing. The control algorithm is validated through simulations on a CSM with parallel and helical actuators’ routing.
5.1.3 Inverse Dynamics
To exploit the mechanical impedance of continuum robots, the controller requires a knowledge of the dynamics of the elastic rod. Furthermore, the dynamic models allow us to consider the influence of external forces, such as gravity or contacts. Therefore, an inverse dynamic model can be used, compensating undesired nonlinear effects and disturbances.
In the case of shape control, Mbakop et al. (2021a) proposed an inverse dynamics controller for a soft finger. The control algorithm is based on a mix of a geometrical and mechanical approach, combining the PH and the Euler-Bernoulli model. While the former allows the reconstruction of the finger’s shape, the latter maps actuators’ magnitude to the positions of PH control points. They experimentally validated the control law with two prototypes: (i) a planar Fluidic Elastomeric Actuator soft finger and a pneumatic CSM. Applying an external tip load to the CSM proved that combining the PH and EBRT improves accuracy and the computational time, w.r.t. using only the PH model.
The subsequent simulation study proposed a new decentralized Inverse Dynamic controller based on the CRT (Doroudchi and Berman, 2021). The controller considers all strain modes and incorporates the system dynamics inside different gain matrices. Each piece or unit of the soft arm is active and independent from the other pieces. This approach increases the computational efficiency and successfully tracks configurations requiring shear deformations with average normalized root-mean-square errors below .
5.1.4 Feedback Linearization
Continuum robots have an inherent under-actuation, which makes control design more difficult. Feedback Linearization (FL) is a widely used method for controlling under-actuated mechanisms (Spong, 1994). The main idea of this control algorithm is to linearize the system with a feedback control action, compensating all the nonlinear terms. The FL controller can be combined with a linear controller (e.g. Linear Quadratic Regulator), in order to achieve the desired performances.
For a task-space control, Rucker et al. (2022) proposed a controller for continuum robots based on a highly under-actuated planar Kirchhoff rod (i.e., the EBRT). The main idea is to show that a maximal coordinate form of the Kirchhoff model is suitable for control purposes, exploiting the under-actuation of the system. The controller is derived by solving a constrained Lagrangian problem, discretizing the length of the Kirchhoff rod. The control algorithm is based on an input-output Feedback Linearization combined with an Sliding Mode Control (SMC) in an outer loop. An observer based on the constrained model formulation provides control feedback, adding corrective virtual external forces based on the estimation error. Lastly, simulations of an actual two-DoFs soft prototype evaluate the control framework and demonstrate the consistency of the algorithms during tip position trajectory tracking. The numerical examples of the SMC also demonstrate robustness to parameter variations.
Caradonna et al. (2024) presented a 2D soft segment with a revolute joint at its base controlled by a FL controller. By analyzing the equilibrium map of the system’s zero dynamics, they proposed a method to employ the system for the pick-and-place task, accounting for the added mass from the object. They utilize the FL controller with the Curvature Parameterization (27) and Strain Parameterization (11) and compare the two discretization techniques emphasizing numerical robustness, computational cost, and accuracy. The control algorithm and pick-and-place strategy validated through simulations demonstrated consistency.
5.1.5 Energy-Shaping
Another well-known control algorithm designed for under-actuated systems is the Energy-Shaping Control (ESC) (Spong, 1996). The main idea is to find an input such that the system reaches a desired energy state. The rod models are particularly effective in this case, thanks to the Hamiltonian formulation of the elastic rods.
Caasenbrood et al. (2021, 2022) proposed an ESC for CSM for tip pose regulation. To apply the control algorithm, the Authors reformulated the GVS model in a Port-Hamiltonian form, showing the passivity condition of the system and enabling the design of the ESC. The control law is validated through numerical examples, proving the capability of the controlled system to reach the desired tip pose. Furthermore, thanks to the properties of the specific control algorithm, both under-actuation and hyper-redundancy are effectively managed. The controlled system demonstrates smooth convergence to the target, and a pseudo-joint strategy similar to that observed in nature.
Chang et al. (2023) introduced an ESC for a CSM inspired by an octopus tentacle employing the CRT with the DER discretization. They modeled three main types of muscles: transverse, longitudinal, and oblique. The transverse muscles necessitate the DER approach due to the invalid assumption of a rigid cross-section. Leveraging the system’s passivity properties, they designed an ESC by solving an optimal control problem based on Hamiltonian principles. Additionally, they developed an efficient method to design a task-specific equilibrium configuration by solving an optimization problem in the group. The model and ESC are validated on the PyElastica simulator, controlling the CSM for reaching and grasping a static object.
Tiwari and Banavar (2023) proposed a ESC for a 2D flexible single-link manipulator, based on CRT. In particular, the model used is geometrically exact, i.e. the manipulator’s configuration space lies on an infinite-dimensional Lie Group. To implement the controller in the discrete-time domain, the Authors used a lie group variational integrator, derived from the Lagrangian variational principle of the flexible single-link manipulator (Tiwari and Banavar, 2023). The results in silico indicate that the discrete closed-loop system achieves shape regulation during point-to-point rotational maneuvers.
5.1.6 Sliding Mode
An alternative to the above controllers for under-actuated systems is the SMC (Xu and Ümit Özgüner, 2008), due to the inherent robustness. The main idea of this strategy is to control the nonlinear system with a discontinuous control signal, forcing it to a specific curve on the phase space (e.g., a line).
An adaptive SMC based on CRT is proposed by Alqumsan et al. (2019). The controller is designed for planar tip position control, which considers the saturation of the actuation sources. Furthermore, the marginal stability of the proposed adaptive control law is provided. The validation is conducted in simulation, where the PDE of the Cosserat rod are solved by the Generalized- method (Chung and Hulbert, 1993), an implicit solver that assures unconditional stability and numerical dissipation for high-frequency noise.
Subsequently, the high nonlinearity of a large-scale and ultralight two-section inflatable CSM made of TPU-coated Nylon is captured by a dynamic Cosserat rod recursively to predict the behavior of the thin-walled and incomplete continuum structure (Li et al., 2021). An extended inverse dynamic model compensates for the error in a feed-forward manner. Moreover, stiffness and damping matrices are calibrated on deflection data and a weight release experiment, respectively. The study also compares the rod model versus a rigid-body and a PCC model. Finally, three closed-loop model-based controllers for trajectory tracking are compared to further demonstrate the viability of the CRT. Here, an adaptive SMC generated the desired pressure differences in the presence of model uncertainties. The chattering phenomenon caused by the conventional SMC is attenuated by integrating a fuzzy logic control, using continuous membership functions to replace the discontinuous .
Mishra et al. (2023) proposed a Terminal SMC (Venkataraman and Gulati, 1993) for tip position TT control. The model on which the controller is based consists of a combination of CRT and fractional Bounc-Wen model (Mishra et al., 2022), taking into account the hysteresis effect. In this work, the Terminal SMC is robust, fast, non-singular, and with exponential convergence. In addition, an obstacle avoidance algorithm is presented, using a combination of cumulative danger field and damped least square methods. The model and the controller are validated in a simulated version of Robotino-XT (Wang et al., 2022a).
5.1.7 Model Predictive Control
The Model Predictive Control (MPC) (Rawlings et al., 2017) is a well-known controller based on iterative and finite-horizon optimization. Furthermore, it is possible to optimize while satisfying constraints, such as actuators’ saturation.
Towards MIS, Soltani et al. (2017) introduce a 3D, stable, non-linear model-based hybrid force/position controller to apply a constant force to the cardiac tissue while tracking the desired trajectory. A quasi-static Cosserat rod models the distal shaft of tendon-driven catheters, reformulated for controller design. The position controller is based on a non-linear MPC tracker satisfying input constraints and local stability conditions. The position control operates in an inner loop, while the force control is in an outer control loop. Simulations conducted with a bidirectional ablation catheter predict the control performance applied in catheterization. Finally, experimental validation is carried out on a real pig heart, though the authors suggest the necessity of more computational efficiency and incorporation of imaging modalities with position tracking systems.
Recently, a novel motion planning method of fish-like soft robots actuated by macro-fiber composite pairs was developed (Barbosa et al., 2023). The structure should mimic oscillatory and undulation movements, accomplished if the amplitude of the tail motion is greater than that of the head motion. The motion planning adopts MPC for generating the actuation signals. This model-based strategy requires a small-sized model. Therefore, they extracted a linear modal state-space model from a free-free dynamic EBRT considering the electro-mechanical coupling of the actuator pairs. Moreover, design strategies like the addition of concentrated and distributed masses were addressed to mimic fish-like motion, guaranteeing suitable mode shapes for the structure. Four numerical case studies demonstrated the capability of the proposal for deriving bounded input signals that generate oscillatory and undulation movements even with disturbances.
5.1.8 Miscellaneous
An open-loop optimal control method to control an octopus-inspired planar arm modeled with CRT was proposed (Wang et al., 2021b). The main goal is to use the optimal control approach to replicate two stereotypical motions: (i) reaching (i.e., bending propagation from the base to the tip) and (ii) fetching (i.e., creation of pseudo-joints). The functional cost of the optimal problem is the elastic potential and actuation energy of the arm, with a terminal cost related to the specific task. An iterative forward-backward algorithm solves the optimization. The proposal validated in silico demonstrated the emergence of propagating waves, as observed in stereotypical strategies.
Conversely, Doroudchi et al. (2018) designed three controllers based on the dynamics of a linear Euler-Bernoulli rod. Thanks to the linearized form of the EBRT, the authors can apply the optimal control theory easily. The novelty here is that the optimal controllers are decentralized, with a modular soft robotic arm where each piece has independent actuation and sensorization. They observe that ill-conditioned system matrices, common in discretized beam models, may hinder control convergence. The iterative algorithm is also very sensitive to the initial condition. Furthermore, a third controller based on the gradient descent method provides a bounded solution for the controller gains and reduces the sensitivity to initial conditions. Finally, simulations validated the approach.
To control aerial continuum manipulation systems, Hashemi et al. (2023) proposed a terminal synergetic controller to accomplish tip position TT. After presenting a novel platform that integrates a quadrotor with a tendon-bent CTR, they designed a two-loop control architecture. In particular, the first loop implements a terminal synergetic controller to ensure the continuum robot’s tip follows the desired trajectory. The second loop’s controller manages the quadrotor’s position and velocity. In addition, this last control algorithm is based on a combination of synergetic control and a hybrid feedback law. For the first time, a synergetic controller is applied in the literature, showing robustness against disturbances and model uncertainties. Finally, they validate the control framework simulating an aerial continuum manipulation system.
Controller | References | Rod Model | Task | Validation | ||||||||||||||||||||
Inverse Kinematics |
Inverse Dynamics |
Feedback Linearization |
Energy-Shaping |
Sliding Mode |
Model Predictive Control |
Open-Loop Optimal Controller | Wang et al. (2021b) | CRT (DER) | Reaching, Fetching | Numerical | ||||||||||||||||||||
Doroudchi et al. (2018) | EBRT | Shape Regulation | Numerical | |||||||||||||||||||||
Synergetic Control | Hashemi et al. (2023) | CRT | Tip Position TT | Numerical |
5.1.9 Shape/State Estimation
As seen in Sec. 5.1.1 and in Fig. 7, a general MB shape controller requires feedback in terms of configuration vector and its time derivative . Depending on the spatial discretization, the vector could be difficult or impossible to measure. To tackle this issue, a general MB control framework requires an Estimator, which maps the measures in the estimated state . For continuum and soft robots, this problem is referred to as Shape/State Estimation or Shape/State Reconstruction. Currently, most Shape Estimation methods use geometrical models or Computer Vision algorithms (e.g., (Bezawada et al., 2022; AlBeladi et al., 2021)). However, incorporating prior knowledge of rod-based models can significantly enhance performance.
In the pioneering work of Trivedi and Rahn (2014), the Authors suggest three approaches for Shape Estimation, based on CRT: (i) mounting load cells at the base, (ii) employing cable encoders, and (iii) mounting an inclinometer at the end of each piece. A maximum tip position error of is observed in simulations. Lastly, they use the OctArm VI (Grissom et al., 2006) to experimentally validate all three approaches. Another study proposes a real-time shape-estimation method based on the force-torque for CTR measured at the tubes’ basis (Donat et al., 2021). It extends a shape estimation algorithm for elastic Kirchhoff rods. They modeled a CTR combining planar PCC segments lying on different equilibrium planes. The approach is evaluated with single and two combined additively manufactured tubes at high frequency, achieving a mean deviation of mm along the tube. Differently, Lilge et al. (2022) develop a model for shape and strain estimation of a continuum cable-driven robot using a Gaussian Process regression, a mathematical tool for estimating continuous-time trajectories in . The idea involves substituting (i) time with arc-length and (ii) kinematic and dynamic laws based on CRT. This method efficiently estimates the robot’s shape using noisy measurements from sensors such as strain gauges, tracking coils, and an external camera. The real-world experiments provide excellent performances, with an average tip pose error of mm and . Finally, Yousefi et al. (2023) propose an optimization-based method that simultaneously estimates the shape and the forces acting on a continuum robot by employing the quasi-static CRT. Magnetic localization determines the position of multiple robot points. The method estimates the robot’s shape and force in a wide range of conditions validated experimentally, with and without knowledge of the contact positions.
5.2 Learning-based Controllers
Learning-based controllers leverage the physics of rod models and utilize Machine Learning (ML) to derive parameterized control policies , which provide and end-to-end mapping between desired poses and sensor measurements to motor commands :
The vector of policy parameters are the weights of ANN trained via Supervised Learning (SL) or RL.
We report learning-based controllers in chronological order to highlight the evolution of the literature.
5.2.1 Supervised Learning
SL is viable to address the soft robotic control problem for simple tasks. It relies on pseudo-random motion data of actuations and corresponding task space to train an inverse model (50).
Initial learning strategies used rod models to rapidly prototype controllers. For instance, Thuruthel et al. (2017) developed dynamic models of CSM using recurrent neural networks and presented a trajectory optimization method for task space control. They validated the controller on a Cosserat rod with PCS parameterization and a pneumatic CSM. Recently, Wang and Rojas (2022) developed a closed-loop controller of continuum robots, employing RNN to approximate the forward and inverse dynamics. They extended the CRT in (Till et al., 2019) to account for spine compression and collect motion data for training their forward models. A non-parametric Gaussian process regression compensates for the error between the real robot and the RNN. A hybrid controller alternates between two stand-alone policies. They validated the control architecture on reaching and tracking tasks using a tendon-driven continuum robot with 3D-printed compressible spines. The results show that combining simulated and experimental data improves the control accuracy.
5.2.2 Reinforcement Learning
RL is one of the main ML paradigms and a versatile framework suited for decision-making and control (Sutton and Barto, 2018). The two fundamental objects of an RL problem are the agent and the environment. In soft robotic control, the agent is a learning-based controller implementing an ANN, whereas the environment is everything outside the agent, comprising the soft robot and other objects. Fig. 8 depicts the agent-environment interaction highlighting the role of ANN and rod models. The agent and the environment interact at discrete time steps . At every interaction step, the agent observes the (possibly partial) state of the environment including sensor readings and task information. The agent acts according to a policy to select actions taking the form of low-level controls like torques or pressures. A policy can be a deterministic mapping from states to actions, . Alternatively, a policy can be stochastic, mapping a state to a distribution over actions . The environment transitions into another state according to the dynamics of the rod model. The agent also perceives a scalar reward signal from the environment, indicating the goodness of the current state specifying the control task. The goal of the agent is to maximize its return, the cumulative sum of rewards:
where is a discount factor. Therefore, the goal of RL is to find an optimal policy that maximizes the expected return
The following deep RL contributions are summarized chronologically in Table 6. Satheeshbabu et al. (2019) applied deep Q-learning (DQL) with experience replay for the quasi-static position control of a pneumatic CSM capable of bending and twisting. An open-loop policy was trained in simulation utilizing a static Cosserat rod. They validate it in simulation and on the physical platform, subject to various external loads. To cope with the high-dimensional state-action space in soft robotics, DQL (Mnih et al., 2013) employs an ANN to approximate the value function of state-action pairs, defined as
A deterministic policy can leverage to select the best action in a state as .
The following work increased the dexterity of the CSM and trained a closed-loop controller for precise quasi-static positioning via deep deterministic policy gradient (DDPG) (Satheeshbabu et al., 2020). Experiments in simulation investigated the robustness of the control policy for effects of loading, reachability, and workspace discontinuity. Finally, the controller was deployed on the prototype. The DDPG algorithm also derived a closed-loop controller for the quasi-static positioning of a pneumatic CSM modeled with the KRT (Uppalapati et al., 2020). Interestingly, the CSM was integrated into a mobile platform including a rigid arm and a sensorized gripper. The system was teleoperated and validated in the agricultural task of picking berries using different maneuvering strategies.
The potential of continuum mechanical models is fully exploited in physical interaction tasks. Naughton et al. (2021) effectively applied several deep RL algorithms to train various control policies. The control tasks included point reaching, trajectory tracking, and maneuvering through structured and unstructured obstacles. The controllers were learned and tested in simulation using a synthetic CSM based on a dynamic CRT. The rod was actuated by applying distributed internal torques, modeled via splines characterized by control points and vanishing values at the rod ends. The work showcased a captivating interaction between the rod and the environment. In particular, the maneuvering task required the free end of the rod to reach a position behind fixed obstacles. The task was solved with a reward function that only considered the distance to the target and a penalty term for the occurrence of numerical errors
Interestingly, the rod learned to navigate and interact with the obstacles without an explicit reward term. Finally, Alessi et al. (2023b) applied proximal policy optimization (PPO) to train a closed-loop position control policy for dynamic trajectory tracking with a pneumatic CSM. The controller was trained in simulation leveraging a dynamic Cosserat rod model of the CSM (Alessi et al., 2023a). After training, simulation tests evaluated the generalization capabilities of the policy to new observations, dynamics, and tasks. The experiments included tracking new trajectories subject to unknown external forces or using different material properties (e.g., ). In addition, the learned controller was transferred without retraining to intercept a moving object. The policy has shown good generalization capabilities for all four tests. Subsequently, Alessi et al. (2024) elaborated on the previous simulated environment to account for the dynamics of a dexterous CSM and its interaction with objects. They developed a learning-based control architecture for pose/force control, which enabled dynamic pushing with CSM in the real world. To mitigate the significant sim-to-real gap in soft robotics, they originally applied domain randomization (DR), an effective sim-to-real technique (Zhao et al., 2020b), emphasizing soft materials.
Controller | References |
Task |
Open loop ANN | Satheeshbabu et al. (2019) | CRT () | DQL | Tip Position TT (way-point) | (None) | ||||||
Closed loop ANN | Satheeshbabu et al. (2020) | CRT () | DDPG | Tip Position TT (way-point) | (None) | ||||||
Uppalapati et al. (2020) | KRT () | DDPG | Tip Position TT (way-point) | (None) | |||||||
Naughton et al. (2021) | CRT () | Several | Tip Pose TT, Obstacle maneuvering | ||||||||
Alessi et al. (2023b) | CRT () | PPO | Tip Position TT, Object interception | ||||||||
Alessi et al. (2024) | CRT () | PPO | Tip Pose/Force for Pushing | (DR) |
5.3 Discussion on Rod-based Controllers
5.3.1 On the Model-based Controllers
The emergence of computationally efficient and accurate models enables researchers to develop model-based controllers that can accurately predict and react to the nonlinear nature of the elastic rods. As shown in Table 5, the controllers based on Inverse Kinematics are the most popular choice, achieving tip trajectory tracking tasks in real-world prototypes. Despite the simplified model, the feedback action is able to reject the unmodelled effects, showing the usability of that strategy.
However, a dynamic model is essential to exploit the properties of continuum robots. Most of the proposed controllers adopt well-known approaches from the under-actuation literature, such as FL, ESC, SMC, or MPC. However, only a few proposed controllers presented an experimental validation. The reasons behind this sim-to-real gap could be the lack of robustness or physical limits of the prototypes. While the Inverse Dynamics, FL, and ESC are sensitive to unmodelled effects and parametric uncertainties, the discontinuous output of the SMC is difficult to implement with the most popular actuation technologies (e.g., tendons, pneumatic). Specifically, the actuators’ internal dynamics may limit the effective bandwidth of the system, hence filtering the discontinuous output of the controller. The MPC is a promising but not widely explored controller in the continuum robots literature.
Another relevant aspect of the analyzed works is the discretization technique employed. Two main aspects eased the control design: (i) the advancement of discretization techniques and (ii) the tunability of the accuracy. The former allows the researchers to design the controller with a finite number of state variables and compute efficiently the robot’s kinematics or dynamics. Selecting appropriate discretization techniques is crucial, particularly for determining the minimum frequency necessary to ensure numerical stability. In addition, the DER technique takes into account also cross-section deformations, especially useful for modeling and controlling muscles in the case of bio-inspired robotic systems (Chang et al., 2023). The latter aspect allows for the choice of accuracy levels and the number of strain modes the model can predict. A common approach simplifies the model and designs a controller that treats unmodeled strains as external disturbances. This strategy can enhance the speed of the control loop and its robustness.
Regarding research gaps, it is possible to note the lack of experimental validation for MB controllers. Furthermore, there are few studies involving external forces due to interactions with environments, particularly because of the complexity involved in modeling and estimating contact and friction forces. Further developments on model-based controllers could benefit from using physics-inspired ANN to lessen the computational burden of complex expressions (e.g., the Coriolis matrix) or use them to compute the EoM of the elastic rod, as already explored in (Lutter and Peters, 2023; Liu et al., 2024). Finally, the literature on rod-based planning (Mishra et al., 2023; Bentley et al., 2023) is sparse, likely due to the current maturity of the controllers. Recently, some planning algorithms using only geometrical approaches have been proposed (Rao et al., 2024).
5.3.2 On the Learning-based Controllers
Learning-based controllers utilize the physics of rod models in various applications, with a clear distinction between SL and RL.
Rod models have a marginal role in SL strategies, which used rod models as proxies for motion data collection under predefined open-loop policies. The motion data were employed to train ANN, which efficiently emulate forward or inverse robot models, and replaced rods to facilitate the preliminary validation of control hypotheses. Although SL was widely used in early soft robotic controllers, it limits the robot tasks to simple reaching and tracking due to the challenges in collecting supervised data for physical interaction tasks. Therefore, this approach could fade away in the future.
Conversely, the role of rod models is central for developing complex deep RL control policies, especially to achieve unstructured physical interaction with CSM. Deep RL agents exploited computationally feasible rod models embedded in simulated environments to learn optimal iteratively. Herein, RL not only replaces the burden of collecting and labeling datasets offline but also unlocks a new range of possibilities not fully explored yet. The literature review highlighted the role of rod models in RL policies. Indeed, as reported in Table 6, control tasks evolved from position control for quasi-static tracking to simulated pose control for obstacle maneuvering and pose/force control for dynamic pushing. For instance, initial policies limited the physical interaction to static payloads without addressing the sim-to-real gap (Satheeshbabu et al., 2020). On the other hand, efforts in developing simulated environments (Naughton et al., 2021) combined with careful reward engineering and sim-to-real techniques (Alessi et al., 2024) facilitated recent advancements in physical interaction with CSM. In summary, rod models combined with recent deep RL methods and sim-to-real transfer techniques (Zhao et al., 2020b) advanced the state of the art of soft robot control.
Future developments could explore model-based RL. Here, the term model-based refers to the agent not only learning a policy but also learning a predictive model of the environment used for planning (Moerland et al., 2023). Alternative research lines can be borrowed from rigid robotics and tailored to the nonlinearities of continuum and soft robots (Ibarz et al., 2021).
6 Conclusion
This paper provided an overview of the modeling and control of continuum and soft robots using rod theories. First, we formulated the four main rod theories used in robotics under a unified mathematical framework. Subsequently, we provided mathematical tools for modeling the actuation in soft robotics. This was followed by a classification and summary of the rod-based models for continuum and soft robots across nine classes of deformations, offering a unique perspective on the evolution of the state of the art, together with a discussion on research trends and gaps that emerged. Furthermore, we reviewed recent control strategies leveraging rod models to perform various tasks, followed by a discussion on the encountered approaches as well as our insights. Through an in-depth exploration of various rod-based models and control strategies, the diverse range of applications of rod theories demonstrates their capabilities to capture the complex mechanical behaviors of continuum and soft robots. Despite the remarkable advancements in incorporating rod theories into soft robot models and controllers, there are still areas requiring further research. For instance, accurate modeling of processes inherent in soft materials, such as hysteresis, is still a challenge. Related to the control, addressing efficient sim-to-real policy transfer in learning-based approaches is essential for real-world robot deployment.
This work was supported by the Future and Emerging Technologies (FET) program of the European Commission, through the PROBOSCIS project (G.A. ID 863212), by the US Office of Naval Research Global under Grant N62909-21-1-2033, and in part by Khalifa University under Awards No. RIG-2023-048, RC1-2018-KUCARS
Appendix A Query for models and controllers
The two Scopus advanced search queries used for the models and controllers are respectively:
TITLE-ABS-KEY (“soft” OR “continuum”) AND TITLE-ABS-KEY (robot* OR “arm” OR “manipulator”) AND TITLE-ABS-KEY (“Cosserat” OR “Kirchhoff” OR “Timoshenko” OR “Euler-Bernoulli” OR “Geometrically exact” OR “rod” OR “beam”) AND (theor* OR model*).
TITLE-ABS-KEY (“soft” OR “continuum”) AND TITLE-ABS-KEY (robot* OR “arm” OR “manipulator”) AND TITLE-ABS-KEY (“Cosserat” OR ”Kirchhoff” OR “Timoshenko” OR “Euler-Bernoulli” OR “Geometrically exact” OR “rod” OR “beam”) AND (control*).
Appendix B Lie Algebra
The Tilde Operator is defined as
(55) For a vector , its application gives an skew-symmetric matrix
(56) -
The Hat operator is defined as
(57) Let be . The Hat operator gives
(58) -
The Vee operator is the inverse of the Hat operator
(59) Consequentially, it is possible to write .
The Adjoint Representations consist in two maps: and . The former is defined as
(60) where . The latter is defined as
(61) Furthermore, it is possible to define also the co-adjoint operator , that is